### v1

# Weight fluctuations in (deep) linear neural networks and a derivation of the inverse-variance flatness relation

Abstract
We investigate the stationary (late-time) training regime of single- and two-layer linear neural networks within the continuum limit of stochastic gradient descent (SGD) for synthetic Gaussian data. In the case of a single-layer network in the weakly oversampled regime, the spectrum of the noise covariance matrix deviates notably from the Hessian, which can be attributed to the broken detailed balance of SGD dynamics. The weight fluctuations are in this case generally anisotropic, but experience an isotropic loss. For a two-layer network, we obtain the stochastic dynamics of the weights in each layer and analyze the associated stationary covariances. We identify the inter-layer coupling as a new source of anisotropy for the weight fluctuations. In contrast to the single-layer case, the weight fluctuations experience an anisotropic loss, the flatness of which is inversely related to the fluctuation variance. We thereby provide an analytical derivation of the recently observed inverse variance-flatness relation in a deep linear network model.
I Introduction
I.0.1 Motivation
Deep neural networks have proven to be formidable tools in the area of machine learning Goodfellow et al. (2016); Aggarwal (2018); Murphy (2022). Their ability to generalize from data often hinges on the dynamics of their training processes, notably the stochastic gradient descent (SGD) method. While the success of these networks in practical applications is well-documented, the underlying principles governing their learning dynamics, especially concerning weight fluctuations and loss landscapes, are still not fully understood. In order to address some of these aspects analytically, we focus here on one- and two-layer linear networks trained on synthetic Gaussian data with noisy labels. These simplified models have been proven to share many aspects of the learning dynamics of full nonlinear networks Saxe et al. (2014, 2018); Arora et al. (2019); Advani et al. (2020); Hastie et al. (2022); Braun et al. (2022). In fact, in the stationary regime at the end of training, a deep nonlinear network typically follows a linearized dynamics (multivariate Ornstein-Uhlenbeck process with correlated noise) Mandt et al. (2017); Jastrzebski et al. (2018); Chaudhari and Soatto (2018). We employ here the continuum description of SGD in terms of a Langevin equation, which conveniently allows us to analyse training dynamics and steady-state properties adopting a statistical physics perspective Engel and Van den Broeck (2001); Bahri et al. (2020); Huang (2021).
One motivation of the present study stems from the recently observed inverse relation between the variance of a weight mode and the flatness of the loss along the direction of that mode Feng and Tu (2021). While this inverse variance-flatness relation (IVFR) has been investigated from various perspectives Feng and Tu (2021); Xiong et al. (2022); Yang et al. (2023); Adhikari et al. (2023), we seek here a first-principles derivation in a neural network model. This requires, as a preparatory step, also a detailed understanding of the various sources of anisotropy of the noise and the fluctuations in a single-layer linear network, which we address in the first part of this study. Complementing recent works on this aspect Kunin et al. (2021); Ziyin et al. (2022), we consider in particular the slightly oversampled regime near the interpolation threshold (i.e., the point where the sample number equals the input dimension) and investigate the dependencies on the parameter count, sample number and label noise. We find that arguably the simplest model showing an IVFR is a two-layer linear network. In fact, the inter-layer coupling of the weights is a crucial ingredient responsible for the loss anisotropy, which therefore does not occur in the case of a single-layer network. However, even for a single-layer network, the weight fluctuations are anisotropic except in the highly oversampled regime. In a physical sense, the IVFR indicates that the effective noise temperature is higher in regions where the loss valleys are narrow, thus effectively biasing the model towards flat regions of the loss. Indeed, wide loss valleys are typically associated with improved generalization ability of the network Hochreiter and Schmidhuber (1997); Jastrzebski et al. (2018); Soudry et al. (2018).
Another motivation of our study stems from the observation that, while the (deterministic) gradient flow dynamics of deep linear networks has been addressed in several studies Saxe et al. (2014); Lampinen and Ganguli (2019); Braun et al. (2022), the corresponding layer-wise stochastic gradient flow dynamics has received less attention so far. A notable exception is the online learning in two-layer nets, which can be assessed analytically in certain regimes Saad and Solla (1995); Riegler and Biehl (1995); Goldt et al. (2019); Arnaboldi et al. (2023). Recently, effects of stochasticity in two-layer linear nets have been analyzed to explain the observed bias towards sparse weight distributions Chen et al. (2023). In the present work, we focus on supervised learning of mini-batches of data and study the associated stochastic dynamics of the weights in each layer within a Langevin approach.
I.0.2 Related work
Continuum modeling of SGD. The modeling of SGD in terms of a stochastic differential equation (SDE) – in the sense that the former emerges from a Euler-Maruyama time-discretization of the latter – is an established approach in the literature Mandt et al. (2015, 2016, 2017); Li et al. (2017); Hu et al. (2017); Jastrzebski et al. (2018); Chaudhari and Soatto (2018); Wojtowytsch (2021a, b). In particular, the steady-state dynamics at the end of training can be described as a multivariate Ornstein-Uhlenbeck process Mandt et al. (2017); Jastrzebski et al. (2018); Ali et al. (2020); Kunin et al. (2021); Li et al. (2022). Despite being formally derived in the limit of infinitely small learning rate (LR)Yaida (2019), the continuum approximation to SGD is typically found to be meaningful also for finite LRs Li et al. (2021) and is has been successfully used to study various aspects of the training of neural networks. The noise in SGD plays an important role for the generalization ability of a model as it directs the weights towards wider minima of the loss, which improves generalization Hochreiter and Schmidhuber (1997); Jastrzebski et al. (2018); Zhang et al. (2018); Smith and Le (2018); Zhu et al. (2019); Nguyen et al. (2019); Xie et al. (2021); Chen et al. (2022); Mori et al. (2022); Wu et al. (2022); Yang et al. (2023); Sclocchi et al. (2023) and helps escaping from saddle points Hu et al. (2017); Daneshmand et al. (2018). A crucial property of the noise is its anisotropy, i.e., its dependence on the direction in parameter space Zhu et al. (2019); Xie et al. (2021); Ziyin et al. (2022); Feng and Tu (2021). The noise is often assumed to be of Gaussian additive form, uncorrelated in time (see, however, Ref. Kühn and Rosenow (2023)) with a covariance proportional to the Hessian Kunin et al. (2021); Wu et al. (2021). This latter approximation, however, becomes exact only in the infinite sample limit, as we will discuss in Section II.2.2 below. Notably, the discrepancy between noise covariance and Hessian implies that detailed balance is broken in the stationary state reached by SGD, giving rise to interesting non-equilibrium behavior such as persistent currents Chaudhari and Soatto (2018); Kunin et al. (2021); Adhikari et al. (2023). In the context of linear regression, effects of parameter-dependent (multiplicative) noise have been addressed in Ali et al. (2020). The relation of the noise covariance matrix to the Hessian and the Fisher matrix is further analyzed in Thomas et al. (2020).
(Deep) Linear nets. Single-layer linear networks often admit closed form analytic solutions which qualitatively explain many properties observed in deep nonlinear networks (see, e.g., Krogh and Hertz (1992); Engel and Van den Broeck (2001); Advani et al. (2020); Mei et al. (2019); Hastie et al. (2022)). In the limit of large width, the neural tangent kernel provides an linear approximation of a nonlinear network around initialization Lee et al. (2019); Jacot et al. (2018). The steady-state dynamics of a trained nonlinear network becomes effectively linear – a fact that, e.g., is utilized in the construction of Bayesian sampling methods Mandt et al. (2017); Chaudhari and Soatto (2018). Introducing depth to a linear network leads to nonlinear gradient descent dynamics of the weights and a generally non-convex loss Saxe et al. (2014); Laurent and von Brecht (2018). Deep linear nets have emerged as important toy models for the study of nonlinear nets, enabling fundamental insights into training dynamics, generalization ability, implicit regularisation, and spectral bias Saxe et al. (2014); Lampinen and Ganguli (2019); Arora et al. (2018); Ji and Telgarsky (2019); Du and Hu (2019); Arora et al. (2019); Gidel et al. (2019); Saxe et al. (2018); Tarmoun et al. (2021); Braun et al. (2022); Chen et al. (2023).
Inverse variance-flatness relation (IVFR). The inverse relation between the magnitude of the steady-state weight fluctuations and the flatness of the loss has been first discussed in Feng and Tu (2021). As emphasized in Xiong et al. (2022), using the formalism of Ao (2004); Kwon et al. (2005); Ao and Yin (2006), the actual effective potential (understood as the inverse of the covariance matrix of the weight distribution) is not the loss, but the product of the inverse of the (generally anisotropic) diffusion term and the loss. This indeed follows also from the classic theory of stochastic processes Risken (1989); Gardiner (2009); Kunin et al. (2021) and has been realized in various machine-learning contexts Mandt et al. (2017); Chaudhari and Soatto (2018). In Ref. Yang et al. (2023), IVFR has been derived within a simple dynamic model of slow and fast degrees of freedom. The stability of IVFR under dropout has been discussed in Zhang et al. (2021).
I.0.3 Outline of the present work
We first consider learning of synthetic data via SGD in a single-layer linear network (Section II). Employing the continuum description of the training dynamics, we show that the spectrum of the noise covariance matrix in the stationary state deviates significantly from the Hessian, which we relate to the broken detailed balance of the SGD dynamics. As a consequence, the weight fluctuations in the underparameterized regime are generally anisotropic, i.e., their amplitude depends on the principle direction.
Proceeding to a two-layer linear network (Section III), we determine a continuum description of the layer-wise training dynamics under SGD. We obtain coupled stochastic differential equations for the weights and determine expressions for the drift and diffusion matrices in the large batch regime. These matrices turn out to be generally rank deficient due to the coupling between the network layers. As a consequence the weight fluctuations in each layer are anisotropic.
We finally apply our results to study the shape of the loss under a perturbation along a principle direction of the weight fluctuation. In the case of a single-layer network, the loss remains isotropic despite the anisotropy of the weight fluctuations. For a two-layer network, we thereby confirm the recently observed Feng and Tu (2021) inverse relation between the amplitude of a weight fluctuation and the width of the loss along the principle direction of that fluctuation. The coupling between the layers of the network is crucial for this result. The present work thus provides a first-principles derivation of the inverse variance-flatness relation in an analytically tractable model of a deep neural network.
II Single-layer linear networks
In this section, we consider the fluctuating dynamics of the weights of an underparameterized single-layer linear network. The network evolves under a mean-square loss and thus solves a linear regression problem. The present linear framework can be matched to a nonlinear network in the stationary state at the end of training (see Appendix A).
II.1 Linear regression
We begin by recapitulating the theory of linear regression that is necessary for the subsequent parts of our study Bishop (2006); Hastie et al. (2009). The training data consists of $P$ samples $\mathbf{x}^{\mu}\in\mathbb{R}^{N}$ ($\mu\in 1,\ldots,P$), the components being i.i.d. Gaussian random variables $x^{\mu}_{i}\sim\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ with mean $\mu_{x}$ and variance $\sigma_{x}^{2}=1/N$  [^1] This variance is chosen in order to obtain quantities of $\mathcal{O}(1)$ after applying random teacher weights as given in Eq. 1.. The targets $y^{\mu}\in\mathbb{R}$ are provided by a linear teacher model:
$y^{\mu}=\mathbf{u}^{T}\mathbf{x}^{\mu}+\epsilon^{\mu},$ (1)
with a fixed random weight vector $\mathbf{u}\in\mathbb{R}^{N}$ with $u_{i}\sim\mathcal{N}(0,\sigma_{u}^{2})$ (we typically take $\sigma_{u}^{2}=1$) and a residual error $\epsilon^{\mu}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ modeled as a i.i.d. Gaussian noise. The error term describes the residual dependence between inputs and targets that can not be captured by a linear relationship. In a slight abuse of terminology, we will also in the present regression context refer to $y$ and $\epsilon$ as ‘label’ and ‘label noise’, respectively Kunin et al. (2021); Ziyin et al. (2022). The complete set of the training data $(\mathbf{x}^{\mu},y^{\mu})_{\mu\in\{1,\ldots,P\}}$ is called batch, while a randomly chosen subset $(\mathbf{x}^{\mu},y^{\mu})_{\mu\in\mathcal{B}}$ of size $S<P$ with =$\mathcal{B}=\{\pi(1),\ldots,\pi(S)\}$ (and $\pi$ being a random permutation of the $P$ elements) is called a mini-batch. The case of statistically independent random labels $y^{\mu}$, which sometimes is useful in order to analyze fluctuations, corresponds to the replacements $\mathbf{u}=0$, $\epsilon^{\mu}\to y^{\mu}$. We define a design matrix $X\in\mathbb{R}^{N\times P}$ with the columns representing the samples $\mathbf{x}^{\mu}$, and, correspondingly, a noise vector $\bm{\varepsilon}=(\epsilon^{1},\epsilon^{2},\ldots,\epsilon^{P})\in\mathbb{R}^{1\times P}$, such that Eq. 1 can be written as
$Y=\mathbf{u}^{T}X+\bm{\varepsilon}\in\mathbb{R}^{1\times P}.$ (2)
The linear network is defined by $y=\mathbf{w}^{T}\mathbf{x}$, with the learnable weights $\mathbf{w}\in\mathbb{R}^{N}$. The full-batch loss takes the form of a least-squares regression model:
$L(\mathbf{w})=\frac{1}{2P}||\mathbf{w}^{T}X-Y||^{2}=\frac{1}{2P}(\mathbf{w}^{T}X-Y)(\mathbf{w}^{T}X-Y)^{T}=\frac{1}{2}\left(\mathbf{w}^{T}\Sigma^{xx}\mathbf{w}-\mathbf{w}^{T}\Sigma^{xy}-\Sigma^{xy,T}\mathbf{w}+YY^{T}/P\right),$ (3)
with the moment matrices
$\Sigma^{xx}=H=\frac{1}{P}XX^{T},\qquad\Sigma^{xy}=\frac{1}{P}XY^{T}.$ (4)
Here, $H\in\mathbb{R}^{N\times N}$ denotes the Hessian, which is identical to $\Sigma^{xx}$ for this model [see Eq. 9 below]. Over- and underparameterized regimes correspond to $N>P$ and $N<P$, respectively. We will typically consider $\mu_{x}=0$; hence $\Sigma^{xx,xy}$ represent covariance matrices and, due to the i.i.d. assumption for $\mathbf{x}^{\mu}$, in limit $P\to\infty$ one has $H_{ii}=\sigma_{x}^{2}$, while $H_{ij}=0$ for $i\neq j$. In the following we make use of the Moore-Penrose pseudoinverse $\mathcal{A}^{+}$ of a matrix $\mathcal{A}$. The pseudoinverse can be expressed as $\mathcal{A}^{+}=(\mathcal{A}^{T}\mathcal{A})^{-1}\mathcal{A}^{T}$ if $\mathcal{A}$ has linearly independent columns, and as $\mathcal{A}^{+}=\mathcal{A}^{T}(\mathcal{A}\mathcal{A}^{T})^{-1}$ if $\mathcal{A}$ has linearly independent rows.
Using the elementary derivative rule $\partial(\mathbf{x}^{T}\mathbf{a})/\partial\mathbf{x}=\mathbf{a}=\partial(\mathbf{a}^{T}\mathbf{x})/\partial\mathbf{x}$, the condition for the minimum of the loss can be written as
$\nabla_{\mathbf{w}}L=\Sigma^{xx}\mathbf{w}^{*}-\Sigma^{xy}=0.$ (5)
In the underparameterized case, the solution readily follows as $\mathbf{w}^{*}_{\text{up}}=(\Sigma^{xx})^{-1}\Sigma^{yx}$. By contrast, in the overparameterized case, $\Sigma^{xx}$ is not full rank, and hence infinitely many solutions can be obtained by adding an arbitrary vector of the nullspace of $\Sigma^{xx}$ to any particular solution $w^{*}$. Taking advantage of this freedom, one thus typically considers the minimum norm solution obtained from the constrained minimization problem $\tilde{L}=||\mathbf{w}||^{2}-(\mathbf{w}^{T}X-Y)\bm{\kappa}$, where $\bm{\kappa}\in\mathbb{R}^{P}$ is a vector of Lagrange multipliers Schaeffer et al. (2023). The solution in the overparameterized case then follows as $\mathbf{w}^{*}_{\text{op}}=X(X^{T}X)^{-1}Y^{T}=(XX^{T})^{+}XY^{T}$, where we used the fact that the so-called gram matrix $X^{T}X\in\mathbb{R}^{P\times P}$ now has full rank and applied the general identity $X(X^{T}X)^{+}=(XX^{T})^{+}X$ of the pseudoinverse. Defining
$J=\frac{1}{P}\sum_{\mu}^{P}\mathbf{x}^{\mu}\epsilon^{\mu}=\frac{1}{P}X\bm{\varepsilon}^{T}\in\mathbb{R}^{N}$ (6)
the minimum-norm solution of the linear regression problem can be compactly written as
$\mathbf{w}^{*}=(\Sigma^{xx})^{+}\Sigma^{yx}=\begin{cases}(XX^{T})^{-1}XY^{T}=\mathbf{u}+H^{-1}J,\qquad&N<P\quad\text{(underpar.)}\\ X(X^{T}X)^{-1}Y=H^{+}H\mathbf{u}+H^{+}J,\qquad&N>P.\quad\text{(overpar.)}\end{cases}$ (7)
The quantity $H^{+}H$ can be understood as a projector onto the orthogonal complement of the nullspace of $H$.
Notably, by using only general properties of the pseudo-inverse [and not Eq. 7], the loss in Eq. 3 can be rewritten as Ziyin et al. (2022)
$L(\mathbf{w})=\frac{1}{2}(\mathbf{w}-\mathbf{u}-H^{+}J)^{T}H(\mathbf{w}-\mathbf{u}-H^{+}J)-\frac{1}{2}J^{T}H^{+}J+\frac{1}{2P}\bm{\varepsilon}\bm{\varepsilon}^{T}.$ (8)
Using the symmetry of $H$ and the elementary relation $H^{+}HH^{+}=H^{+}$ of the pseudoinverse, one can readily show that Eq. 8 is minimized by Eq. 7. Accordingly, Eq. 8 can be reformulated as
$L(\mathbf{w})=\frac{1}{2}(\mathbf{w}-\mathbf{w}^{*})^{T}H(\mathbf{w}-\mathbf{w}^{*})-\frac{1}{2}J^{T}H^{+}J+\frac{1}{2P}\bm{\varepsilon}\bm{\varepsilon}^{T},$ (9)
making the minimizing property of $\mathbf{w}^{*}$ apparent. In the overparameterized case, the last two terms in Eq. 9 cancel, such that $L(\mathbf{w}^{*})=0$, showing that the model is able to perfectly fit the data generated via Eq. 1. This cancellation can be proven by using the SVD $X=VD^{1/2}U^{T}$ ($V\in\mathbb{R}^{N\times N}$, $D^{1/2}\in\mathbb{R}^{N\times P}$, $U\in\mathbb{R}^{P\times P}$, with $V$, $U$ being orthogonal and $D^{1/2}$ having $\min(P,N)$ singular values on the diagonal) and the relation $X^{+}=U(D^{1/2})^{+}V^{T}$, which implies that $X^{T}(XX^{T})^{+}X=\mathbbm{1}_{P\times P}$. By contrast, the latter relation does not hold in the underparameterized case, since $U\mathbbm{1}_{P\times P}^{N}U^{T}\neq\mathbbm{1}_{P\times P}$ if $N<P$, where $\mathbbm{1}_{P\times P}^{N}$ is the unit matrix with the last $P-N$ elements on the diagonal being zero.
II.2 Dynamics
Next, we consider the continuum approximation of the stochastic gradient descent, i.e., the stochastic gradient flow (SGF), which we will use to study the spectrum of the noise and its dependence on the number of parameters and on the label noise. We consider the stationary state at the end of training, where the weights fluctuate around the solution $\mathbf{w}^{*}$.
II.2.1 SGD and continuum approximation
With the mini-batch loss ($S=|\mathcal{B}|$ is the mini-batch size)
$L_{B}(\mathbf{w})=\frac{1}{S}\sum_{\mu\in\mathcal{B}}\ell^{\mu}(\mathbf{w}),\qquad\ell^{\mu}(\mathbf{w})=\frac{1}{2}(\mathbf{w}^{T}\mathbf{x}^{\mu}-y^{\mu})^{2},$ (10)
and the learning rate $\lambda$, the weights evolve according to SGD as
$\mathbf{w}(\tau+1)=\mathbf{w}(\tau)-\lambda\nabla_{w}L_{B}(\mathbf{w}(\tau))=\mathbf{w}(\tau)-\lambda\nabla L(\mathbf{w}(\tau))-\lambda\,\bm{\xi}(\mathbf{w}(\tau)),$ (11)
where $\tau$ denotes the discrete time step. The mini-batch noise $\bm{\xi}$ is defined as
$\bm{\xi}(\mathbf{w})\equiv\nabla L_{B}(\mathbf{w})-\nabla L(\mathbf{w})$ (12)
and fulfils $\nabla L=\langle\nabla L_{B}\rangle_{\mathcal{B}}$, where $\langle\cdot\rangle_{B}$ denotes the expectation over the full batch.
In order to proceed to the continuum limit, we follow Kunin et al. (2021); Li et al. (2017); Jastrzebski et al. (2018); Chaudhari and Soatto (2018) and assume the learning rate $\lambda$ to be sufficiently small, such that Eq. 11 can be approximated, in the weak sense, by a SDE (Langevin equation):
$\partial_{t}\mathbf{w}(t)=-\nabla L(\mathbf{w}(t))+\sqrt{\lambda}\,\bm{\xi}(t),\qquad\nabla L(\mathbf{w})=H(\mathbf{w}-\mathbf{w}^{*}),\qquad\langle\bm{\xi}(t)\bm{\xi}(t^{\prime})^{T}\rangle_{\mathcal{B}}=C(\mathbf{w})\delta(t-t^{\prime})$ (13)
where $t=\tau/\lambda$ is a new time variable. The noise $\bm{\xi}(t)$ has weight-dependent Gaussian correlations in weight space, as characterized by the correlation matrix $C(\mathbf{w})$, which is further analyzed below. The noise is interpreted in an Ito sense, since, according to Eq. 11, $\mathbf{w}(\tau)$ is independent of $\mathbf{w}(\tau+1)$ Gardiner (2009). Although the SGD-SDE mapping is not strictly rigorous Yaida (2019); Li et al. (2021), we find that the results obtained from Eqs. 13 and 11 agree well even for rather large learning rates.
II.2.2 Noise covariance
According to Eq. 12, the covariance matrix of the noise, averaged over batches (with replacement), is given by (see Appendix C)
$C=\frac{1}{S}\left[\frac{1}{P}\sum_{\mu=1}^{P}(\nabla\ell^{\mu})(\nabla\ell^{\mu})^{T}-(\nabla L)(\nabla L)^{T}\right].$ (14)
The expression in the case of batching without replacement takes a similar form and approaches the above when $S\ll P$ Ziyin et al. (2022); Wu et al. (2022). In the case of linear regression, we have $\nabla_{w}\ell^{\mu}=(\mathbf{w}^{T}\mathbf{x}^{\mu}-y^{\mu})\mathbf{x}^{\mu}$ [see Eq. 10]. Using Eq. 1, the covariance takes the form:
$C_{ab}=\frac{1}{S}\left[\frac{1}{P}\sum_{\mu}^{P}[\bar{\mathbf{w}}^{T}\mathbf{x}^{\mu}-\epsilon^{\mu}]^{2}x^{\mu}_{a}x^{\mu}_{b}-H\bar{\mathbf{w}}\bar{\mathbf{w}}^{T}H\right],\qquad\bar{\mathbf{w}}\equiv\mathbf{w}-\mathbf{u}.$ (15)
This result shows that, beside the external source of randomness stemming from the mini-batching of input data $\mathbf{x}^{\mu}$ and the label noise $\epsilon^{\mu}$, also the fluctuations of the weights $\mathbf{w}$ contribute to the SGD-noise. Estimating the typical magnitudes of the various terms in Eq. 15 (see Appendix D) indicates that the $w$-dependent terms dominate if $||\bar{\mathbf{w}}^{2}||\sigma_{x}^{2}\gg\sigma_{\epsilon}^{2}$. In the absence of label noise ($\sigma_{\epsilon}^{2}=0$), this condition is fulfilled at all times. By contrast, if $\sigma_{\epsilon}^{2}>0$, we write $\mathbf{w}=\mathbf{w}^{*}+\delta\mathbf{w}$, where $\delta\mathbf{w}$ denotes the weight fluctuations around the stationary solution $\mathbf{w}^{*}$ [Eq. 7]. Since the typical stationary variance $\langle\delta w_{i}^{2}\rangle\sim(\lambda/S)\sigma_{\epsilon}^{2}$ [c.f. Eq. 34 below and Appendix D] and $||H^{-1}J||^{2}\sim\sigma_{\epsilon}^{2}/\sigma_{x}^{2}$, multiplicative noise terms will dominate in the stationary state only if $(\lambda/S)\sigma_{x}^{2}\gg 1$. In the case of non-vanishing label noise ($\sigma_{\epsilon}^{2}>0$), the additive noise approximation is therefore warranted for typical choices of model parameters and normalized data. Exceptions can arise at early stages of the training, where $||\bar{\mathbf{w}}||$ can be large. In the following, we will study in more detail the various approximations to the SGD noise.
Figure 1: (a) Variance of the matrix elements of $\mathcal{P}=X^{+}X$, which enters in Eq. 18, as a function of sample number $P$. Curves for various $N$ and $P$ superimpose when expressed in appropriately scaled coordinates. (b) Entries, ordered by magnitude, of the diagonal matrix $K$ [Eq. 18], which determines the covariance of the SGD noise in Eq. 17. The various curves correspond to sample numbers ranging between $P=70$ and $P=10000$ (from bottom to top; with fixed input dimension $N=50$). The black curve represents the approximation $K^{\mu\mu}=(\bm{\epsilon}^{\mu})^{2}$, where $\bm{\epsilon}^{\mu}$ is a vector of independent Gaussian random numbers of zero mean and variance $\sigma_{\epsilon}^{2}$
Figure 2: (a) Eigenvalues of the noise covariance matrix $C$, ordered by their magnitude, for varying number of samples (increasing from bottom to top) in the overparameterized regime for $N=100$ and varying sample numbers. Numerical results obtained from SGD (solid lines) are compared to the theoretical predictions obtained from Eq. 17. (b) Eigenvalues of $C$ given by Eq. 17 (solid lines) compared to those of Eq. 20 (broken lines, essentially corresponding to the eigenvalues of the Hessian $H$), for various $P/N=50,2,1.1$ (from top to bottom) at $N=100$. The inset shows the maximum and minimum eigenvalues of $C$ from Eq. 17 (solid lines) and Eq. 20 (dashed lines) for varying $P$. Eigenvalues in (a,b) are normalized by the factor $\sigma_{x}^{2}\sigma_{\epsilon}^{2}/S$, where $\sigma_{x}^{2}=1/N$ is the variance of the samples [see also Eq. 20] (c) Visualization of the noise covariance (for $N/P\simeq 0.67$) in the basis of the Hessian $H$ [Eq. 4], i.e., $\tilde{C}=\langle|V^{T}CV|\rangle_{x,\epsilon}$, where the average is over several data realizations and the matrix $V$ consists of the eigenvectors of $H$ as columns. The color range is scaled logarithmically and normalized to the maximum entry of $\tilde{C}$. The magnitude of the off-diagonal entries decrease upon increasing $P$.
Additive noise approximation for finite sample number ($P<\infty$). The additive noise approximation consists of evaluating $C$ [Eq. 15 on the stationary solution $\mathbf{w}=\mathbf{w}^{*}$ [Eq. 7], assuming low noise magnitude ($S\gg 1$). In the overparameterized case, each individual loss gradient vanishes for $\mathbf{w}=\mathbf{w}^{*}$, i.e., $\nabla_{\mathbf{w}}\ell^{\mu}=0$. By contrast, in the underparameterized case, only the sum over the total batch is zero: $\sum_{\mu}^{P}\nabla_{\mathbf{w}}\ell^{\mu}=0$. Consequently, the noise and the weight fluctuations vanish in the overparameterized regime in the steady-state [see also the discussion after Eq. 9]:
$C=0\quad\text{if $N>P$}.$ (16)
This also applies for output dimensions larger than one, i.e., $\mathbf{y}^{\mu}\in\mathbb{R}^{d_{o}}$ with $d_{o}>1$, because the different output channels are not coupled by the loss.
Since generally $J\neq 0$ [see Eq. 6] for finite sample number, Eq. 15 takes the form
$C_{ab}=\frac{1}{S}\left[\frac{1}{P}\sum_{\mu}^{P}(J^{T}H^{+}\mathbf{x}^{\mu}-\epsilon^{\mu})^{2}x^{\mu}_{a}x^{\mu}_{b}\right]=\frac{1}{S}(XKX^{T})_{ab},$ (17)
with a diagonal matrix in sample space:
$K^{\mu\nu}=\frac{1}{P}\left[\sum_{j,k}^{N}\sum_{\gamma}^{P}\frac{1}{P}\epsilon^{\gamma}x_{k}^{\gamma}H^{+}_{kj}x^{\mu}_{j}-\epsilon^{\mu}\right]^{2}\delta^{\mu\nu}=\frac{1}{P}\left[\left(\frac{1}{P}\bm{\varepsilon}X^{T}H^{+}X-\bm{\varepsilon}\right)^{\mu}\right]^{2}\delta^{\mu\nu}.$ (18)
Note that these expressions are general and do not rely on simplifying assumptions on the data or the label noise. In the case of fully random labels $y^{\mu}$, they apply with the replacement $\epsilon^{\mu}\to y^{\mu}$. Similar forms of $C$ in various limits have also been obtained in Ali et al. (2020); Chen et al. (2021).
In the overparameterized case, one readily shows, using identities for the pseudoinverse, that $X^{T}(XX^{T})^{+}X=\mathbbm{1}$, hence $K=0$. In the underparameterized case, one has $X^{T}(XX^{T})^{+}X=X^{+}X=:\mathcal{P}$, which is a projector on the row space of $X$ and is independent of the variance of $X$. In the large sample regime $P\gg N$, a numerical analysis (see Fig. 1) reveals that the variance of the elements of this matrix behaves as $\mathrm{var}(\mathcal{P})\propto P^{-1.8}N^{-1}$ (the variance is slightly larger on the diagonal), while the mean value along the diagonal $\langle\mathcal{P}_{ii}\rangle\simeq N/P$. For the purpose of estimating the variance of $\varepsilon\mathcal{P}$, we write $\mathcal{P}\simeq(N/P)\mathbbm{1}_{P}+P^{-0.9}N^{-0.5}\tilde{\mathcal{P}}$, where $\tilde{\mathcal{P}}$ is a random matrix with i.i.d. entries of zero mean and unit variance. Accordingly, we infer $\mathrm{var}(\bm{\varepsilon}X^{+}X)_{ij}\sim P\sigma_{\epsilon}^{2}\mathrm{var}(\mathcal{P})\to 0$ as $P\to\infty$, which implies $K^{\mu\nu}\simeq(\bm{\varepsilon}^{\mu})^{2}\delta^{\mu\nu}$, i.e., in the relevant oversampled regime and for $P\sigma_{\epsilon}^{2}\gg 1$, $K$ essentially becomes a diagonal matrix of squared Gaussian random numbers. This is illustrated in Fig. 1, where the diagonal entries of $K$ (ordered by magnitude) are displayed for various sample sizes $P$.
In Fig. 2, the spectrum of the noise covariance matrix obtained from numerical simulations in the stationary state is shown and compared to the theoretical prediction of Eq. 17 averaged over the distributions of the data $x$ and the label noise $\epsilon$. We generally find good agreement with the additive noise approximation, which we have checked for various label noise strengths $\sigma_{\epsilon}^{2}$ and learning rates $\lambda$. As shown in Fig. 2(b), the noise spectrum significantly deviates from the Hessian approximation of Eq. 20, but approaches the latter in the limit $P/N\to\infty$. The deviation is caused by the weak cross-correlations between the data eigenmodes induced by the label noise [Eq. 17], as illustrated in Fig. 2(c). This becomes also clear by noting that, from the SVD of the design matrix $X=VD^{1/2}U^{T}$ (with $V\in\mathbb{R}^{N\times N}$, $U\in\mathbb{R}^{P\times N}$, and a positive-semidefinite diagonal matrix $D^{1/2}\in\mathbb{R}^{N\times N}$), the noise covariance in the Hessian eigenbasis takes the form $V^{T}CV=\frac{1}{S}D^{1/2}U^{T}KUD^{1/2}$, which differs from $D/S$ since $U^{T}KU\neq\mathbbm{1}$.
Large sample number limit ($P\to\infty$). In this regime, using the Gaussian i.i.d. assumption of the data and label noise, $C$ can be evaluated as Ziyin et al. (2022)  [^2] It is useful to note that $(\nabla L)(\nabla L)^{T}=H\bar{\mathbf{w}}\bar{\mathbf{w}}^{T}H$, which cancels one such term stemming from the first term in Eq. 14.
$C=\frac{1}{S}\left[H\bar{\mathbf{w}}\bar{\mathbf{w}}^{T}H+\mathrm{tr}(H\bar{\mathbf{w}}\bar{\mathbf{w}}^{T})H+\sigma_{\epsilon}^{2}H\right],\qquad P\to\infty,$ (19)
which explicitly reveals the multiplicative character of the noise. Using the fact that $J=0$ [see Eq. 6] for $P\to\infty$ (since $x^{\mu}$ and $\epsilon^{\mu}$ are uncorrelated), we have $\mathbf{w}\simeq\mathbf{w}^{*}=\mathbf{u}$ in the stationary state [see Eq. 7]. Setting thus $\bar{\mathbf{w}}=0$ in Eq. 19 renders
$C\simeq\frac{\sigma_{\epsilon}^{2}}{S}H,\qquad P\to\infty,$ (20)
which constitutes an often made approximation of the SGD noise covariance in terms of the Hessian Mori et al. (2022); Kunin et al. (2021); Jastrzebski et al. (2018).
Multiplicative noise process for $\sigma_{\epsilon}^{2}=0$. Returning to the general case $\mathbf{w}\neq\mathbf{w}^{*}$, further insight into the nature of the stochastic process associated with SGD can be gained in the case of vanishing label noise $\varepsilon=0$. In this situation, Eq. 14 becomes
$C_{ab}=\frac{1}{S}\left[\frac{1}{P}\sum_{\mu}^{P}\bar{w}_{i}x_{i}^{\mu}x_{a}^{\mu}x_{b}^{\mu}x_{j}^{\mu}\bar{w}_{j}-\left(\frac{1}{P}\bar{w}_{i}\sum_{\mu}x_{i}^{\mu}x_{a}^{\mu}\right)\left(\frac{1}{P}\sum_{\mu}x_{b}^{\mu}x_{j}^{\mu}\bar{w}_{j}\right)\right]=\frac{1}{S}\mathrm{cov}_{x}[(\bar{\mathbf{w}}\cdot\mathbf{x})\mathbf{x}]_{ab}.$ (21)
Accordingly, Eq. 13 can be formulated as
$\partial_{t}\bar{\mathbf{w}}=-H\bar{\mathbf{w}}+\sqrt{\frac{\lambda}{S}}\bar{\mathbf{w}}^{T}Z,$ (22)
with a matrix-valued Gaussian white noise $Z_{ij}$ with covariance ($\Delta x_{i}x_{j}\equiv x_{i}x_{j}-\langle x_{i}x_{j}\rangle$)
$\langle Z_{ij}(t)Z_{kl}(t^{\prime})\rangle=\langle(\Delta x_{i}x_{j})(\Delta x_{k}x_{l})\rangle\delta(t-t^{\prime})=(\langle x_{i}x_{k}\rangle\langle x_{j}x_{l}\rangle+\langle x_{i}x_{l}\rangle\langle x_{j}x_{k}\rangle)\delta(t-t^{\prime}).$ (23)
In the infinite sample limit ($P\to\infty$), the i.i.d. assumption (as well as $\mu_{x}=0$) for the $x_{i}$ implies $\langle Z_{ij}(t)Z_{kl}(t)\rangle=(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk})\langle x^{2}\rangle\delta(t-t^{\prime})$ or, equivalently, $C_{ab}=S^{-1}(\bar{w}_{a}\bar{w}_{b}+\bar{\mathbf{w}}^{2}\delta_{ab})\langle x^{2}\rangle^{2}$. The noise thus generally couples the different weights $\bar{w}_{i}$ during the evolution of Eq. 22, even for whitened data (where $H=\mathbbm{1}_{N}$). In the case of $N=1$ dimensions (but arbitrary finite $P$), the noise covariance simplifies to $\langle(\Delta x^{2})^{2}\rangle=\langle x^{4}\rangle-\langle x^{2}\rangle^{2}=2\langle x^{2}\rangle^{2}$ and Eq. 22 reduces to the standard geometric Brownian motion process Ali et al. (2020)
$\partial_{t}\bar{w}=-\langle x^{2}\rangle\bar{w}+\sqrt{\frac{2\lambda}{S}}\langle x^{2}\rangle\bar{w}\zeta,\qquad\langle\zeta(t)\zeta(t^{\prime})=\delta(t-t^{\prime}),$ (24)
where $\zeta$ is a Gaussian white noise. The mean and variance of $\bar{w}$ for this process are given by $\langle\bar{w}(t)\rangle=\bar{w}(0)e^{-\langle x^{2}\rangle t}$ and $\langle\bar{w}(t)^{2}\rangle=\bar{w}(0)^{2}\left[e^{2\langle x^{2}\rangle\left(\frac{\lambda}{S}\langle x^{2}\rangle-1\right)t}-e^{-2\langle x^{2}\rangle t}\right]$, showing that both converge to zero provided $\langle x^{2}\rangle<S/\lambda$ (which is fulfilled for typical parameter choices). If $\langle x^{2}\rangle=S/\lambda$ and $\bar{w}(0)^{2}=\lambda\langle x^{2}\rangle/S=1$, mean and variance become identical to those of an Ornstein-Uhlenbeck process.
II.2.3 Relaxational dynamics
Figure 3: Relaxation of the loss $L(\mathbf{z})$ (dash-dotted curve) and the weight modes $z_{k}(t)=V^{T}_{kj}w_{j}$ [for $k=0,10,20$, see Eq. 25] towards the solution $z_{k}^{*}$. Modes obtained from SGD (solid) are compared to theoretical predictions $\langle z_{k}(t)\rangle$ (Eq. 27, averaged over the noise; dashed), for a slightly underparameterized linear net with $P/N\simeq 1.04$. The inset shows the loss and the mode $|z_{0}(t)|$ in double-logarithmic representation. The gray line represents the theoretical prediction $L^{*}$ for the final loss.
In order to study the dynamics induced by the SGD, we project the weights onto the eigenmodes of the data. Assuming the SVD of the design matrix to be $X=P^{1/2}VD^{1/2}U^{T}$, with $V\in\mathbb{R}^{N\times N}$, $U\in\mathbb{R}^{P\times N}$, and a positive-semidefinite diagonal matrix $D^{1/2}\in\mathbb{R}^{N\times N}$ (the sample number $P$ is introduced in order for the entries of $D$ to coincide with the eigenvalues of $H$), the Langevin Eq. 13 can be transformed to
$\dot{\mathbf{z}}=-D(\mathbf{z}-\mathbf{z}^{*})+\sqrt{\lambda}\bm{\eta},\qquad\langle\bm{\eta}\bm{\eta}^{T}\rangle=\frac{P}{S}D^{1/2}U^{T}KUD^{1/2},$ (25)
where $\mathbf{z}\equiv V^{T}\mathbf{w}\in\mathbb{R}^{N}$ is the new dynamical variable, $\mathbf{z}^{*}\equiv V^{T}\mathbf{w}^{*}$ are the transformed solution weights [see Eq. 7], and $\bm{\eta}=V^{T}\bm{\xi}$ is the new noise. Analogously, the loss [Eq. 9] can be expressed as
$L=\frac{1}{2}(\mathbf{z}-\mathbf{z}^{*})^{T}D(\mathbf{z}-\mathbf{z}^{*})+L^{*}=\frac{1}{2}\sum_{k}^{N}(z_{k}-z_{k}^{*})^{2}\omega_{k}+L^{*},$ (26)
with $L^{*}\equiv L(\mathbf{z}^{*})=-\frac{1}{2}J^{T}H^{+}J+\frac{1}{2P}\bm{\varepsilon}\bm{\varepsilon}^{T}=\frac{1}{2P}\bm{\varepsilon}(\mathbbm{1}_{P}-X^{T}(XX^{T})^{-1}X)\bm{\varepsilon}^{T}$ denoting the loss of the fully trained model. Equation 25 describes a multivariate Ornstein-Uhlenbeck process driven by correlated Gaussian noise. Owing to the presence of label noise, encoded in the matrix $K$ [see Eq. 18], the covariance of the noise $\bm{\eta}$ is not diagonal but instead induces correlations between the weight modes $z_{i}$. Denoting the relaxation rates as $\omega_{k}\equiv D_{kk}$, ordered such that $\omega_{0}\leq\omega_{1}\leq\ldots$, the solution of Eq. 25 follows as
$z_{k}(t)=z_{k}^{*}+e^{-\omega_{k}(t-t_{0})}[z_{k}(t_{0})-z_{k}^{*}]+\int_{t_{0}}^{t}\mathrm{d}s\,e^{\omega_{k}(s-t)}\eta_{k}(s).$ (27)
For times larger than the smallest relaxation time ($t\gg 1/\omega_{0}$), the influence of the initial value term $\mathbf{z}(t_{0})-\mathbf{z}^{*}$ has decayed and $\mathbf{z}$ fluctuates around the equilibrium solution $\mathbf{z}^{*}$. Near the interpolation threshold $N=P$, the eigenvalues of the Hessian $H$ become arbitrarily small, i.e., $\omega_{0}\to 0^{+}$ Advani et al. (2020); Rocks and Mehta (2020); Singh et al. (2022). Specifically, for Gaussian data we consider here [see Section II.1], the Marchenko-Pastur law of random matrix theory implies $\omega_{0}=\sigma_{x}^{2}(\sqrt{P/N}-1)^{2}$ in the regime $P,N\to\infty$ with $P/N=\mathrm{const}$ Bun et al. (2017); Advani et al. (2020)  [^3] In the overparameterized case ($P<N$), additionally a number of exactly zero eigenvalues appears..
The typical relaxation behavior of the modes near the interpolation threshold is illustrated in Fig. 3. While for $P\gg N$ the relaxation behavior is accurately described by Eq. 27, we find that, for $P\simeq N$, modes $z_{k}(t)$ for small $k$ notably deviate from the purely exponential relaxation, showing sign changes and two-step relaxation. An interesting observation (see inset to Fig. 3) is furthermore that the loss appears to settle to the constant $L^{*}$ already for $t\ll\omega_{0}$, although the model has not yet reached the stationary state. This is a consequence of the presence of the factor $\omega_{k}$ in the loss [Eq. 26], due to which the contribution of modes pertaining to the smallest $\omega_{k}$ can be hidden below the noise floor $L^{*}$, especially for large label noise $\sigma_{\epsilon}^{2}$.
II.3 Steady-state properties
In the following, we analyze the stationary covariance matrix of the weights induced by the SGF of a single-layer linear network. We focus, in particular, on the spectral properties of the weight fluctuations and their dependence on the sample number and the label noise. We then use these results to show that the IVFR is generally not fulfilled for a single-layer linear net.
II.3.1 Steady-state solution
Figure 4: Statistics of weight fluctuations for a linear network in the stationary state. (a) Behavior of the weight variance as a function of input dimension for $P=200$ samples. Simulation results (connected symbols) obtained for SGD with $\lambda=0.1$, $\sigma_{\epsilon}^{2}=0.25$ and $S=10$ are compared to the theoretical predictions of Eqs. 34 and 30 for $\langle\langle\delta w_{i}(t)^{2}\rangle_{t}\rangle_{i}=N^{-1}\sum_{i}M_{ii}$ (dotted and dash-dotted lines), which represents the temporal variance of the weights in the stationary state [where $\mathbf{w}(t\to\infty)=\mathbf{w}^{*}$, see Eq. 7] averaged over all weights. The quantity $\langle\delta w_{i}(t\to\infty)^{2}\rangle_{i}$ (filled circles and dashed line) represents the variance across all weights at an arbitrary but fixed time in the stationary state. The train and test losses are also shown for comparison. (b) Eigenvalues of the covariance matrix of the weights $M=\langle\delta\mathbf{w}\delta\mathbf{w}^{T}\rangle$, ordered by their magnitude and normalized by $\lambda\sigma_{\epsilon}^{2}/(2S)$ [see Eq. 34], for varying number of samples ($P/N=1.05,1.1,1.2,1.5,2.0,4.0$, from bottom to top) in the underparameterized regime (for input dimension $N=100$ and learning rate $\lambda=0.1$). Numerical results obtained from SGD (solid lines) are compared to the theoretical predictions obtained from Eq. 30 (dashed lines). The straight (dashed-dotted) black line represents the prediction of Eq. 34, while the dotted lines represent the theoretical eigenvalues under the detailed balance assumption [Eq. 33]. (c) Smallest and largest eigenvalues of $M$ as given by Eq. 30 (solid lines) or Eq. 33 (dot-dashed lines), and of $Q$ [Eq. 30, dashed lines], normalized by $\lambda\sigma_{\epsilon}^{2}/(2S)$ and for varying sample numbers $P$. The inset shows the dependence on $p$ of the ratio of the Frobenius norms of $Q$ and $C$ (solid line), and of $M$ from Eq. 30 and $M$ from Eq. 34 (dashed lines).
The Fokker-Planck equation for the distribution $p(t,\mathbf{w})$ induced by Eq. 13 takes the form Risken (1989); Gardiner (2009):
$\partial_{t}p(t,\mathbf{w})=-\sum_{i}\frac{\partial}{\partial w_{i}}J_{i},\qquad J_{i}=-\left[\frac{\partial}{\partial w_{i}}L(\mathbf{w})\right]p(t,\mathbf{w})-\sum_{j}\frac{\partial}{\partial w_{j}}\cdot\left[\frac{\lambda}{2}C_{ij}(\mathbf{w})p(t,\mathbf{w})\right],$ (28)
with $L(\mathbf{w})=H(\mathbf{w}-\mathbf{w}^{*})$. We focus in the following on the $w$-independent noise defined by $C$ in Eq. 17, which, as discussed above, is expected to apply in the stationary state. The steady-state covariance matrix of the weights $M\equiv\langle\delta\mathbf{w}\delta\mathbf{w}^{T}\rangle$ (with $\delta\mathbf{w}\equiv\mathbf{w}-\mathbf{w}^{*}$) is determined by the Lyapunov equation Risken (1989); Laub (2005); Mandt et al. (2017); Kunin et al. (2021)
$HM+MH^{T}=\lambda C.$ (29)
The most general solution of this equation fulfills $\nabla\cdot\mathbf{J}=0$, but not necessarily $\mathbf{J}=0$, i.e., while the probability density remains stationary [see Eq. 28], nonzero currents are still permitted (broken detailed balance). This situation is indeed realized for SGD and one obtains Chaudhari and Soatto (2018); Kunin et al. (2021):
$M$ $\equiv\langle\delta\mathbf{w}\delta\mathbf{w}^{T}\rangle=\frac{\lambda}{2}H^{-1}(C+Q),$ (30)
$Q$ $=V\tilde{Q}V^{T},\qquad\tilde{Q}_{ij}=\frac{\omega_{i}-\omega_{j}}{\omega_{i}+\omega_{j}}(V^{T}CV)_{ij},$ (31)
where $\omega_{i}$ are the eigenvalues and the columns of $V$ contain the eigenvectors of $H$ so that $H=V\mathrm{diag}(\omega_{i})V^{T}$. The stationary probability distribution and current follow as
$p_{\text{ss}}(\mathbf{w})=\frac{1}{Z}\exp\left[-\frac{1}{2}(\mathbf{w}-\mathbf{w}^{*})^{T}M^{-1}(\mathbf{w}-\mathbf{w}^{*})\right],\qquad\mathbf{J}_{\text{ss}}=-QM^{-1}(\mathbf{w}-\mathbf{w}^{*})p_{\text{ss}}(\mathbf{w}),$ (32)
with a normalization factor $Z$. The inverse covariance matrix $M^{-1}$ can be interpreted as an effective potential $U(\mathbf{w})\equiv(1/2)\delta\mathbf{w}^{T}M^{-1}\delta\mathbf{w}$, which, due to the presence of the kinetic term $(C+Q)^{-1}$, differs in general from the loss Chaudhari and Soatto (2018); Xiong et al. (2022). The matrix $Q$ is anti-symmetric ($Q=-Q^{T}$) and determines a stationary current $\mathbf{J}_{\text{ss}}$, which flows in planes of constant force, i.e., $\mathbf{J}_{\text{ss}}\cdot(\nabla U)=0$ Kwon et al. (2005); Kunin et al. (2021). In order to aid comparison between various conventions used, e.g., in numerical codes, we note that, if the loss [Eq. 10] is scaled by a factor $\alpha$, i.e., $\ell^{\mu}\to\alpha\ell^{\mu}$, the noise covariance [Eq. 14] scales as $C\to\alpha^{2}C$ (analogously $Q\to\alpha^{2}Q$) and the Hessian $H\to\alpha H$, which implies $M\to\alpha M$.
A solution fulfilling detailed balance can be obtained from straightforward integration of the condition $\mathbf{J}=0$ in Eq. 28, which yields $\frac{1}{2}\lambda C_{ij}\cdot\partial\ln p/\partial w_{j}=-\partial L(\mathbf{w})/\partial w_{i}$ and eventually
[…]
While, generally, detailed balance is broken for SGD, we find that Eq. 33 constitutes a reasonable approximation sufficiently deep in the underparameterized regime (specifically, for $P/N\gtrsim 2$, see discussion of Fig. 4 below). In fact, asymptotically for $P/N\to\infty$, one has $C\propto H$ [see Eq. 20], which implies $V^{T}CV\propto\delta_{ij}$ and thus $Q=0$ in Eq. 30. In this limit detailed balance is thus exactly fulfilled and one obtains isotropic weight fluctuations Kunin et al. (2021); Ziyin et al. (2022),
$M=\frac{\lambda\sigma_{\epsilon}^{2}}{2S}\mathbbm{1},\qquad P/N\to\infty.$ (34)
In Fig. 4, the behavior of the weight fluctuations of a fully trained linear network is illustrated. As shown in panel (a), the steady-state variance $\langle\langle\delta w_{i}(t)^{2}\rangle_{t}\rangle_{i}$ obtained from SGD simulations (by computing the temporal variance of each weight $w_{i}$ and the averaging over all weights $i=1,\ldots,N$) agrees with the theoretical prediction given by the diagonal entries of $M$ in Eq. 30. Notably, the variance is non-vanishing only for $N/P<1$, since an overparameterized network can fit the data in Eq. 1 exactly [see also the comment after Eq. 15], whereas, in the limit $N/P\to 0$, the variance approaches the isotropic result in Eq. 34 (dotted horizontal line). For comparison, the train and test loss are included in the plot, where the latter shows a “double descent” peak at the interpolation threshold ($P=N$), characteristic for learning data with noisy labels Opper et al. (1990); Krogh and Hertz (1992); Belkin et al. (2019); Advani et al. (2020); Hastie et al. (2022); Nakkiran et al. (2019). Beside the temporal variance, we show also the spatial variance of the weights $\langle\delta w_{i}(t\to\infty)^{2}\rangle_{i}=\langle\delta(w_{i}^{*})^{2}\rangle$, which peaks at $P=N$ Chen et al. (2021). We recall that a stationary state is only reached for non-vanishing label noise. By contrast, for $\sigma_{\epsilon}=0$, the network can fit the teacher weights exactly such that the loss exponentially approaches zero (up to machine precision).
Figure 4(b) illustrates eigenvalues of the weight covariance matrix $M=\langle\delta\mathbf{w}\delta\mathbf{w}^{T}\rangle$, ordered by their magnitude, for various sample numbers $P$ in the underparameterized regime. The numerical results from SGD are well captured by Eq. 30 (dashed lines), while strong deviations from Eq. 34 are noticeable, indicating broken detailed balance except for large $P$. In Fig. 4(c), the smallest and largest eigenvalues of the weight covariance matrix $M$ [Eqs. 30 and 34, solid and dot-dashed lines, respectively] are shown as functions of the reduced sample number $P/N-1$. We infer that the detailed balance approximation for $M$ given in Eq. 33 is applicable for $P\gtrsim 2N$. This is also corroborated in the inset, where the Forbenius norms of $Q$ [Eq. 31] and $C$ [Eq. 17] are compared. From the approach of the minimum and maximum eigenvalues we finally conclude that the weight fluctuations become essentially isotropic [see Eq. 34] for $P\gtrsim 100N$, consistent with the behavior of the noise eigenvalues in Fig. 2(b). Before proceeding, we summarize some important steady-state properties for linear regression in the following Table 1.
student weights [Eq. 7]
final loss [Eq. 9]
temporal weight fluctuations
underparameterized
approach teacher weights as $\epsilon\to 0$
nonzero in general; vanishes $\sim\sigma_{\epsilon}^{2}$
nonzero in general; vanish $\sim\sigma_{\epsilon}^{2}$ [Eq. 34]
overparameterized
differ from teacher weights
vanishes (also if $\sigma_{\epsilon}^{2}>0$)
vanish [Eqs. 16 and 20]
Table 1: Properties of a linear network in the stationary (late-time) regime, trained on a teacher-student task [see Eq. 1]. Note that these properties are independent of the sampling method (i.e., sampling with or without replacement). The variance of the target (label) noise is denoted by $\sigma_{\epsilon}^{2}$.
II.3.2 Loss perturbation
Figure 5: Perturbing the linear regression loss [Eq. 9] along eigenvectors of the empirical weight covariance matrix $M=\langle\delta\mathbf{w}\delta\mathbf{w}^{T}\rangle$ obtained from SGD (learning rate $\lambda=0.1$) of a linear network. The insets show the eigenvalues of the weight covariance matrix (ordered by their magnitude vs. mode index $i$). (a) Perturbed loss curves $\delta L(\theta\mathbf{v}_{i})$ for various eigenvectors $\mathbf{v}_{i}$ for $N=100$ dimensional input and $P=2000$ Gaussian i.i.d. samples. The simulation results superimpose on the theoretical prediction $\delta L=\frac{1}{2}\theta^{2}\sigma_{x}^{2}$. (b) Perturbed loss $\delta L(\theta\mathbf{v}_{i})/\theta^{2}$ at fixed (arbitrary) $\theta$ as a function of eigenmode index $i$ for $P=200$ Gaussian correlated samples with $N=40$ dimensions. The black dashed line represents the theoretical prediction for i.i.d. data. In the correlated case, $\delta L$ shows strong random fluctuations as a function of mode index, which decrease upon averaging over multiple realizations of the data.
We now investigate whether the IVFR Feng and Tu (2021) applies in the context of a single-layer linear network. Perturbing the linear regession loss in Eq. 9 around the solution $\mathbf{w}^{*}$ [Eq. 7] renders
$\delta L(\theta\mathbf{v})=L(\mathbf{w}^{*}+\theta\mathbf{v})-L(\mathbf{w}^{*})=\frac{\theta^{2}}{2}\mathbf{v}^{T}H\mathbf{v},$ (35)
where $\mathbf{v}$ is assumed to be an eigenvector of the weight covariance matrix $M$ [Eq. 30]. Owing to the different structural forms of $M$ and $H$, we generally do not expect $\delta L(\theta\mathbf{v})$ to depend in any particular way on the covariance eigenvalues, even if the latter significantly vary in mangitude. This is confirmed by SGD simulations for Gaussian independent and Gaussian correlated data (see Fig. 5), as well as for purely random labels $y^{\mu}$.
Analytic insight can be obtained in the large (infinite) sample number regime, where Eq. 34 implies the eigenvectors $\mathbf{v}\in\{\mathbf{e}_{i}\}_{i=1,\ldots,N}$ with a common eigenvalue $\gamma=\lambda\sigma^{2}/S$. This yields $\delta L(\theta\mathbf{e}_{i})=\frac{1}{2}\theta^{2}H_{ii}=\frac{1}{2}\theta^{2}\langle x_{i}^{2}\rangle$, i.e., the loss perturbation reveals the variance of the $i$th feature. Accordingly, unless one engineers a particular variance dependence on the feature index, the loss perturbation behaves in an essentially random and isotropic way. We thus conclude that the IVFR is not fulfilled for a typical single-layer linear network.
III Two-layer linear network
We now transfer the Langevin formalism presented in the preceding section to a two-layer linear network, which will enable us to analyze the fluctuations of the weights in the individual layers. Since the stochastic dynamics of the weight matrices is now nonlinear, we develop a linearization around the deterministic solution. We will finally use these results in order to derive the IVFR.
III.1 Model
We consider training data consisting of $P$ samples $\mathbf{x}^{\mu}\in\mathbb{R}^{N_{i}}$, $\mathbf{y}^{\mu}\in\mathbb{R}^{N_{o}}$ ($\mu=1,\ldots,P$), with i.i.d. Gaussian $x_{i}^{\mu}\sim\mathcal{N}(0,\sigma_{x}^{2})$ and $y_{i}^{\mu}\sim\mathcal{N}(0,\sigma_{y}^{2})$, which, for simplicity, have zero mean. The targets (‘labels’) $\mathbf{y}^{\mu}$ can, in principle, be created from a linear transformation of the $\mathbf{x}^{\mu}$ as in Eq. 1, although we do not suppose such a relation in the following. The data is conveniently organized in terms of the design matrices $X\in\mathbb{R}^{N_{i}\times P}$, $Y\in\mathbb{R}^{N_{o}\times P}$. The (full-batch) loss of a linear 2-layer network is given by
$L(W_{1},W_{2})$ $=\frac{1}{2P}\sum_{\mu=1}^{P}||\mathbf{y}^{\mu}-W_{2}W_{1}\mathbf{x}^{\mu}||^{2}=\frac{1}{2P}\mathrm{tr}\left[(Y-wX)(Y-wX)^{T}\right]$ (36)
$=\frac{1}{2}\mathrm{tr}\left(\Sigma_{0}^{xx}w^{T}w-w^{T}\Sigma_{0}^{yx}-\Sigma_{0}^{yx,T}w+YY^{T}/P\right),$ (37)
where $W_{1}\in\mathbb{R}^{N_{h}\times N_{i}}$, $W_{2}\in\mathbb{R}^{N_{o}\times N_{h}}$ are the weight matrices, $w\equiv W_{2}W_{1}\in\mathbb{R}^{N_{o}\times N_{i}}$ denotes the product matrix, and
$\Sigma_{0}^{xx}=\frac{1}{P}\mathbf{X}\mathbf{X}^{T},\qquad\Sigma_{0}^{yx}=\frac{1}{P}\mathbf{Y}\mathbf{X}^{T}$ (38)
are the full-batch input-input and output-input covariance matrices. The minimum of the loss is determined by the linear regression solution
$w^{*}\equiv(W_{2}W_{1})^{*}=\Sigma_{0}^{yx}(\Sigma_{0}^{xx})^{+},$ (39)
which effectively fixes only $N_{\text{eff}}=N_{o}N_{i}$ degrees of freedom out of the total number of parameters $N_{\text{tot}}=N_{h}(N_{i}+N_{o})$. As was the case for a single-layer net [see Eq. 16], a deep linear network of the form considered here has non-vanishing weight fluctuations in the stationary state provided $N_{i}<P$, irrespective of the output dimension $N_{o}$. We will thus restrict ourselves to this regime in the following, in which case the pseudo-inverse in Eq. 39 accordingly becomes the usual inverse. In certain cases, where indicated, we will also assume $N_{i}\geq N_{h}\geq N_{o}$.
Introducing the mini-batch covariance matrices as
$\Sigma^{xx}(t)=\frac{1}{S}\sum_{\mu\in\mathcal{B}(t)}\mathbf{x}^{\mu}\mathbf{x}^{\mu,T},\qquad\Sigma^{yx}(t)=\frac{1}{S}\sum_{\mu\in\mathcal{B}(t)}\mathbf{y}^{\mu}\mathbf{x}^{\mu,T},$ (40)
where $\mathcal{B}$ denotes the set of sample indices in the mini-batch (of size $S=|\mathcal{B}|$), the SGD equations take the form
$\frac{1}{\lambda}\left[W_{1}(\tau+1)-W_{1}(\tau)\right]$ $=-\frac{\partial L(w,\tau)}{\partial W_{1}}=W_{2}^{T}\left(\Sigma^{yx}-W_{2}W_{1}\Sigma^{xx}\right),$ (41a)
$\frac{1}{\lambda}\left[W_{2}(\tau+1)-W_{2}(\tau)\right]$ $=-\frac{\partial L(w,\tau)}{\partial W_{2}}=\left(\Sigma^{yx}-W_{2}W_{1}\Sigma^{xx}\right)W^{T}_{1},$ (41b)
with the learning rate $\lambda$.
The full-batch dynamics of Section III.1 in the limit of small $\lambda$ (gradient flow) has been analyzed previously Saxe et al. (2014, 2018); Braun et al. (2022); Gunasekar et al. (2017). We will derive the continuum limit of Section III.1 in the next section, but already note here that, in the case of gradient flow with an explicitly time-dependent loss $L$, one can show that the following quantity is conserved for a linear network ($t=\tau/\lambda$):
$W_{2}^{T}(t)W_{2}(t)-W_{1}(t)W_{1}^{T}(t)=\mathrm{const}.$ (42)
The proof of this relation follows the same steps as in Arora et al. (2018), where it was demonstrated for linear networks with losses that do not explicitly depend on time. When the weights are initialized such that the constant matrix on the r.h.s. of Eq. 42 vanishes, the weights are conventionally called zero-balanced.
Notably, for whitened inputs $\Sigma_{0}^{xx}\propto\mathbbm{1}$ and zero-balanced weights, an explicit solution of this gradient flow can be constructed Braun et al. (2022)  [^4] The proof presented in Appendix C of Braun et al. (2022) should apply for arbitrary $N_{i}$ and $N_{o}$.: Denoting the compact singular value decomposition of $\Sigma_{0}^{yx}$ by
$\Sigma_{0}^{yx}=\tilde{U}\tilde{S}\tilde{V}^{T},\qquad\text{with}\qquad\tilde{U}\in\mathbb{R}^{N_{o}\times N_{m}},\tilde{S}\in\mathbb{R}^{N_{m}\times N_{m}},\tilde{V}\in\mathbb{R}^{N_{i}\times N_{m}},$ (43)
where $\tilde{S}$ is a diagonal matrix with $N_{m}$ nonzero eigenvalues on its diagonal, the late time solution can be expressed as
[…]
However, explicit expressions for the individual $W_{i}$ are not available. Unless otherwise mentioned, we will in the following not assume any special initial conditions.
III.2 Continuum description of SGD
We now derive the continuum limit of the SGD dynamics in each layer, which renders two coupled nonlinear SDEs. A crucial step is to determine the covariances of the effective noise terms induced by the mini-batching. In order to analyze the nonlinear SDEs in the stationary state, we focus on the large batch-regime, where fluctuations around the deterministic gradient flow solution are small.
III.2.1 Noise sources and stochastic gradient flow
Since the only source of randomness enters via the mini-batching of $\Sigma^{xx,yx}$ Han et al. (2021), we have
$\Sigma^{xx}=\Sigma^{xx}_{0}+\delta\Sigma^{xx},\qquad\Sigma^{yx}=\Sigma^{yx}_{0}+\delta\Sigma^{yx},$ (45)
where we note that $\Sigma_{0}^{xx}=\langle\Sigma^{xx}\rangle_{\mathcal{B}}$ and $\Sigma_{0}^{yx}=\langle\Sigma^{yx}\rangle_{\mathcal{B}}$. Accordingly, the noise terms are defined as
$\delta\Sigma^{xx}=\Sigma^{xx}-\Sigma_{0}^{xx},\qquad\delta\Sigma^{yx}=\Sigma^{yx}-\Sigma_{0}^{yx},$ (46)
and their covariances can be expressed (using Appendix C) by a four-dimensional tensor:
$\begin{split}\mathrm{cov}(\delta\Sigma^{xx})_{ij,kl}\equiv\langle\delta\Sigma^{xx}_{ij}\delta\Sigma^{xx}_{kl}\rangle_{\mathcal{B}}&=\left\langle\left(\frac{1}{S}\sum_{\mu}x^{\mu}_{i}x^{\mu}_{j}-\Sigma_{0,ij}^{xx}\right)\left(\frac{1}{S}\sum_{\nu}x^{\nu}_{k}x^{\nu}_{l}-\Sigma_{0,kl}^{xx}\right)\right\rangle_{\mathcal{B}}\\ &=\frac{1}{S}\left[\frac{1}{P}\sum_{\mu=1}^{P}x^{\mu}_{i}x^{\mu}_{j}x^{\mu}_{k}x^{\mu}_{l}-\Sigma_{0,ij}^{xx}\Sigma_{0,kl}^{xx}\right].\end{split}$ (47)
In general, a four-dimensional tensor $\mathcal{M}_{ijkl}$ can be reshaped into a conventional two-dimensional matrix $\mathpzc{m}$ by defining $\mathpzc{m}_{ab}=[\mathcal{M}_{ij,kl}]_{a=(j-1)N_{i}+i,b=(l-1)N_{k}+k}$ where $N_{i}$ and $N_{k}$ are the sizes of dimensions indexed by $i$ and $k$, respectively. This definition is consistent with the usual vectorization $\mathpzc{m}=\mathrm{vec}(\mathcal{A})\mathrm{vec}(\mathcal{B})^{T}$ in the case where $\mathcal{M}_{ijkl}=\mathcal{A}_{ij}\mathcal{B}_{kl}$, which applies to the covariance tensor. The vectorization operation $\mathrm{vec}(\mathcal{A})$ is defined by stacking the columns of the matrix $\mathcal{A}$ on top of each other to form a column vector.
In order to arrive at tractable expressions, we evaluate the noise covariance in the large sample number limit ($P\to\infty$), which is analogous to the Hessian approximation given by Eq. 20. Practically, the Hessian approximation for the noise applies already for $P\gtrsim 100N$, as discussed in Section II.3.1. In this limit, the first term in the square brackets in Eq. 47 converges to a batch average, such that Wick’s rule for products of Gaussian variables can be applied: $\frac{1}{P}\sum_{\mu=1}^{P}x^{\mu}_{i}x^{\mu}_{j}x^{\mu}_{k}x^{\mu}_{l}\simeq\langle x_{i}x_{j}x_{k}x_{l}\rangle_{\mathcal{B}}=\langle x_{i}x_{j}\rangle_{\mathcal{B}}\langle x_{k}x_{l}\rangle_{\mathcal{B}}+\langle x_{i}x_{k}\rangle_{\mathcal{B}}\langle x_{j}x_{l}\rangle_{\mathcal{B}}+\langle x_{i}x_{l}\rangle_{\mathcal{B}}\langle x_{j}x_{k}\rangle_{\mathcal{B}}$. Hence, altogether one obtains
$\mathrm{cov}(\delta\Sigma^{xx})_{ij,kl}$ $=\frac{1}{S}\left[\Sigma_{0,ik}^{xx}\Sigma_{0,jl}^{xx}+\Sigma_{0,il}^{xx}\Sigma_{0,jk}^{xx}\right],$ (48a)
$\mathrm{cov}(\delta\Sigma^{yx})_{ij,kl}$ $=\frac{1}{S}\left[\langle y_{i}^{2}\rangle\delta_{ik}\Sigma_{0,jl}^{xx}+\Sigma_{0,il}^{yx}\Sigma_{0,kj}^{yx}\right],$ (48b)
$\mathrm{cov}(\delta\Sigma^{xx}_{ij},\delta\Sigma^{yx}_{kl})$ $=\frac{1}{S}\left[\Sigma^{yx}_{0,ki}\Sigma_{0,jl}^{xx}+\Sigma_{0,il}^{xx}\Sigma_{0,kj}^{yx}\right],$ (48c)
where we used the factor that the $y_{i}$ are uncorrelated zero mean random variables. In the case of whitened inputs, i.e.
$\Sigma_{0}^{xx}=\sigma_{x}^{2}\mathbbm{1},$ (49)
the above covariances simplify to
$\mathrm{cov}(\delta\Sigma^{xx})_{ij,kl}=\frac{1}{S}\left[\sigma_{x}^{4}\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}\right],\qquad\mathrm{cov}(\delta\Sigma^{yx})_{ij,kl}=\frac{1}{S}\left[\langle y_{i}^{2}\rangle\sigma_{x}^{2}\delta_{ik}\delta_{jl}+\Sigma_{0,il}^{yx}\Sigma_{0,jk}^{yx}\right].$ (50)
Upon inserting these expansions of the data covariance matrices into Section III.1 and proceeding to the continuum limit analogously to the single-layer case [Eq. 13], we obtain the following stochastic gradient flow (SGF) equations:
$\partial_{t}W_{1}$ $=W_{2}^{T}\left(\Sigma_{0}^{yx}-W_{2}W_{1}\Sigma_{0}^{xx}\right)+\mathcal{R}_{1},$ (51a)
$\partial_{t}W_{2}$ $=\left(\Sigma_{0}^{yx}-W_{2}W_{1}\Sigma_{0}^{xx}\right)W^{T}_{1}+\mathcal{R}_{2},$ (51b)
with the multiplicative noise matrices $\mathcal{R}_{1}\in\mathbb{R}^{N_{h}\times N_{i}}$, $\mathcal{R}_{2}\in\mathbb{R}^{N_{o}\times N_{h}}$:
$\mathcal{R}_{1}$ $=\sqrt{\lambda}\,W_{2}^{T}\big{(}\delta\Sigma^{yx}-W_{2}W_{1}\delta\Sigma^{xx}\big{)},$ (52a)
$\mathcal{R}_{2}$ $=\sqrt{\lambda}\big{(}\delta\Sigma^{yx}-W_{2}W_{1}\delta\Sigma^{xx}\big{)}W_{1}^{T}.$ (52b)
III.2.2 Perturbation expansion
In order to analyze the stationary fluctuation dynamics, we introduce a suitable linearization into Section III.2.1. Accordingly, we decompose the weight matrices as $W_{1,2}=W_{1,2}^{0}+\delta W_{1,2}$, where $W_{1,2}^{0}$ denotes the solution of the full-batch GD and $\delta W_{1,2}$ describe small fluctuations around this solution. Upon expanding Section III.2.1 up to linear order in the fluctuations of $W$ and $\Sigma$, and collecting terms of the same order, we obtain the deterministic gradient flow equations
$\partial_{t}W_{1}^{0}$ $=W^{0,T}_{2}\left(\Sigma_{0}^{yx}-W_{2}^{0}W_{1}^{0}\Sigma_{0}^{xx}\right),$ (53a)
$\partial_{t}W_{2}^{0}$ $=\left(\Sigma_{0}^{yx}-W_{2}^{0}W_{1}^{0}\Sigma_{0}^{xx}\right)W_{1}^{0,T},$ (53b)
and the dynamics of the first order perturbations:
$\partial_{t}\delta W_{1}$ $=\big{(}\delta W_{2}^{T}\Sigma_{0}^{yx}-\delta W^{T}_{2}w^{*}\Sigma_{0}^{xx}-W^{0,T}_{2}\delta W_{2}W_{1}^{0}\Sigma_{0}^{xx}-W_{2}^{0,T}W_{2}^{0}\delta W_{1}\Sigma_{0}^{xx}\big{)}+\mathcal{R}_{1},$ (54a)
$\partial_{t}\delta W_{2}$ $=\big{(}\Sigma_{0}^{yx}\delta W_{1}^{T}-\delta W_{2}W_{1}^{0}\Sigma_{0}^{xx}W_{1}^{0,T}-W_{2}^{0}\delta W_{1}\Sigma_{0}^{xx}W_{1}^{0,T}-w^{*}\Sigma_{0}^{xx}\delta W_{1}^{T}\big{)}+\mathcal{R}_{2}.$ (54b)
in which the noises [Section III.2.1] now become additive (we assume the noises to be quantities of the same order as $\delta W_{1,2}$):
$\mathcal{R}_{1}$ $=\sqrt{\lambda}\,W_{2}^{0,T}\big{(}\delta\Sigma^{yx}-W_{2}^{0}W_{1}^{0}\delta\Sigma^{xx}\big{)}=\sqrt{\lambda}\,W_{2}^{0,T}\big{(}\delta\Sigma^{yx}-w^{*}\delta\Sigma^{xx}\big{)},$ (55a)
$\mathcal{R}_{2}$ $=\sqrt{\lambda}\big{(}\delta\Sigma^{yx}-W_{2}^{0}W_{1}^{0}\delta\Sigma^{xx}\big{)}W_{1}^{0,T}=\sqrt{\lambda}\big{(}\delta\Sigma^{yx}-w^{*}\delta\Sigma^{xx}\big{)}W_{1}^{0,T}.$ (55b)
We focus henceforth on the late-time, (quasi-)stationary state of training, where the $W_{1,2}^{0}$ are essentially time independent or at least change on a significantly longer time scale than the $\delta W_{1,2}$, allowing us to analyse Sections III.2.2 and III.2.2 with $W_{1,2}^{0}\simeq\mathrm{const}$. Since explicit expressions for $W_{i}^{0}$ in terms of the data and initial conditions are not available, we will consider the $W_{1,2}^{0}$ as given outcomes of the gradient flow [Section III.2.2], obtained, e.g., from a numerical solution  [^5] Note that the zero-balanced solution in Eq. 44 involves different matrix products of the $W_{i}$ from the ones in Section III.2.3 and thus cannot be directly used here..
For whitened inputs [see Eq. 49], we have $w^{*}=\Sigma_{0}^{yx}/\sigma_{x}^{2}$ [see Eq. 7] as well as
$(\Sigma_{0}^{yx}\Sigma_{0}^{yx,T})_{ij}=\frac{1}{P^{2}}\sum_{\mu,\nu}\sum_{k}y_{i}^{\mu}x_{k}^{\mu}x_{k}^{\nu}y_{j}^{\nu}\simeq\frac{1}{P^{2}}\sum_{\mu}(y^{\mu})^{2}\sigma_{x}^{2}\delta_{ij}=\frac{1}{P}\sigma_{x}^{2}\langle y^{2}\rangle\delta_{ij},$ (56)
where we used $\mathbf{x}^{\mu}\cdot\mathbf{x}^{\nu}=\sigma_{x}^{2}\delta^{\mu\nu}$ and the fact that $\mathbf{y}^{\mu}$ are i.i.d. variates. With these simplifications, the noise covariance takes the following form (see Appendix E):
$\mathrm{cov}(\mathcal{R}_{1})_{ij,kl}$ $=\sigma_{\mathcal{R}}^{2}\delta_{jl}\left(1-1/P\right)(W_{2}^{0,T}W_{2}^{0})_{ik},\qquad j,l=1,\ldots,N_{i},$ (57a)
$\mathrm{cov}(\mathcal{R}_{2})_{ij,kl}$ $=\sigma_{\mathcal{R}}^{2}\delta_{ik}\left(1-1/P\right)(W_{1}^{0}W_{1}^{0,T})_{jl},\qquad i,k=1,\ldots,N_{o},$ (57b)
$\mathrm{cov}(\mathcal{R}_{1},\mathcal{R}_{2})_{ij,kl}$ $=\sigma_{\mathcal{R}}^{2}(1-1/P)W^{0,T}_{2,ik}W^{0,T}_{1,jl}\qquad\text{with}\qquad\sigma_{\mathcal{R}}^{2}\equiv\frac{1}{S}\lambda\sigma_{x}^{2}\langle y^{2}\rangle,$ (57c)
while Section III.2.2 becomes
$\partial_{t}\delta W_{1}$ $=-\sigma_{x}^{2}\big{[}W_{2}^{0,T}W_{2}^{0}\delta W_{1}+W^{0,T}_{2}\delta W_{2}W_{1}^{0}\big{]}+\mathcal{R}_{1}=-\frac{\partial\tilde{L}}{\partial\delta W_{1}}+\mathcal{R}_{1},$ (58a)
$\partial_{t}\delta W_{2}$ $=-\sigma_{x}^{2}\big{[}\delta W_{2}W_{1}^{0}W_{1}^{0,T}+W_{2}^{0}\delta W_{1}W_{1}^{0,T}\big{]}+\mathcal{R}_{2}=-\frac{\partial\tilde{L}}{\partial\delta W_{2}}+\mathcal{R}_{2},$ (58b)
with the reduced loss $\tilde{L}(\delta W_{1},\delta W_{2})\equiv\frac{1}{2}\sigma_{x}^{2}\,\mathrm{tr}\left[W_{2}^{0,T}W_{2}^{0}\delta W_{1}\delta W_{1}^{T}+\delta W_{2}^{0,T}W_{2}^{0}\delta W_{1}W_{1}^{0,T}+\delta W_{2}^{T}\delta W_{2}W_{1}^{0}W_{1}^{0,T}\right]$ quadratic in the fluctuations.
Recalling the matrix rank equation $\mathrm{rank}(AB)=\min(\mathrm{rank}A,\mathrm{rank}B)$, we note that, for one-dimensional outputs ($N_{o}=1$), one has $\mathrm{rank}(W_{2}^{0})=1=\mathrm{rank}(W_{2}^{0,T}W_{2}^{0})$ and thus $\mathrm{rank}([\mathrm{cov}(\mathcal{R}_{1})])=N_{i}$ according to Eq. 57a. For zero-balanced initial conditions [i.e. Eq. 42 with vanishing r.h.s.], one has $\text{rank}([\mathrm{cov}(\mathcal{R}_{2})])=\text{rank}(W_{1}^{0}W_{1}^{0,T})=\text{rank}(W_{2}^{0,T}W_{2}^{0})=1$, while for random initial conditions, typically $\text{rank}([\mathrm{cov}(\mathcal{R}_{2})])=\text{rank}(W_{1}^{0}W_{1}^{0,T})=N_{h}$.
III.2.3 Transformations
In order to further analyze Section III.2.2, we expand $\delta W_{1,2}$ in the eigenbasis of $W_{1,2}^{0}$. To this end, we introduce the thin SVD $W_{1}^{0}=AS_{1}B^{T}$ [where $A$, $S_{1}$ and $B$ are $(N_{h},N_{h}),(N_{h},N_{i})$ and $(N_{i},N_{i})$ matrices], and the standard SVD $W_{2}^{0}=US_{2}V^{T}$ [where $U$, $S_{2}$ and $V$ are $(N_{o},N_{o}),(N_{o},N_{o})$ and $(N_{h},N_{o})$ matrices], which yield
$W_{1}^{0}W_{1}^{0,T}$ $=AD_{1}A^{T}\quad\text{with}\quad D_{1}\equiv S_{1}S_{1}^{T}\in\mathbb{R}^{N_{h}\times N_{h}},$ (59a)
$W_{2}^{0,T}W_{2}^{0}$ $=VD_{2}V^{T}\quad\text{with}\quad D_{2}\equiv S_{2}^{T}S_{2}\in\mathbb{R}^{N_{o}\times N_{o}}.$ (59b)
The eigenvalues of $W_{1}^{0}W_{1}^{0,T}$ and $W_{2}^{0,T}W_{2}^{0}$ encoded in the diagonal positive-semidefinite matrices $D_{1,2}$ will play a crucial role in the following. Introducing new dynamical variables $z_{1}\equiv V^{T}\delta W_{1}\in\mathbb{R}^{N_{o}\times N_{i}}$, $z_{2}\equiv\delta W_{2}A\in\mathbb{R}^{N_{o}\times N_{h}}$, and noises $\eta_{1}\equiv V^{T}\mathcal{R}_{1}\in\mathbb{R}^{N_{o}\times N_{i}}$, $\eta_{2}\equiv\mathcal{R}_{2}A\in\mathbb{R}^{N_{o}\times N_{h}}$, reduces Section III.2.2 to
$\dot{z}_{1}$ $=-D_{2}z_{1}-(US_{2})^{T}z_{2}S_{1}B^{T}+\eta_{1},$ (60a)
$\dot{z}_{2}$ $=-z_{2}D_{1}-US_{2}\,z_{1}BS_{1}^{T}+\eta_{2}.$ (60b)
The covariance of the new noises $\eta_{1,2}$ follows readily by using the SVDs in Section III.2.2:
$\mathrm{cov}(\eta_{1})_{mj,nl}$ $=\sigma_{\mathcal{R}}^{2}\delta_{jl}D_{2,mn},\qquad(j,l=1,\ldots N_{i},\quad m,n=1,\ldots,N_{o}),$ (61a)
$\mathrm{cov}(\eta_{2})_{im,kn}$ $=\sigma_{\mathcal{R}}^{2}\delta_{ik}D_{1,mn},\qquad(m,n=1,\ldots N_{h},\quad i,k=1,\ldots,N_{o}),$ (61b)
$\mathrm{cov}(\eta_{1},\eta_{2})_{mj,kn}=\mathrm{cov}(\eta_{2},\eta_{1})_{nk,jm}$ $=\sigma_{\mathcal{R}}^{2}(US_{2})_{mk}^{T}(BS_{1}^{T})_{jn},\qquad(j=1,\ldots,N_{i},\quad n=1,\ldots,N_{h},\quad m,k=1,\ldots,N_{o}).$ (61c)
Since the $D_{i}$ are diagonal, these results represent a significant simplification over Sections III.2.2 and III.2.2.
We now introduce the joint vector $\mathbf{Z}\equiv\{\mathrm{vec}(z_{1}),\mathrm{vec}(z_{2})\}\in\mathbb{R}^{N_{o}(N_{i}+N_{h})}$, $\bm{\eta}\equiv\{\mathrm{vec}(\eta_{1}),\mathrm{vec}(\eta_{2})\}$, which allows us to express Section III.2.3 in the compact form
$\dot{\mathbf{Z}}=-\Gamma\mathbf{Z}+\bm{\eta}.$ (62)
The drift matrix $\Gamma$ and the diffusion matrix $\Delta\equiv\langle\bm{\eta}\bm{\eta}^{T}\rangle$ turn out to be identical symmetric square matrices of size $[N_{o}(N_{i}+N_{h})]^{2}$:
$\Gamma/\sigma_{x}^{2}=\begin{pmatrix}\mathbbm{1}_{N_{i}}\otimes D_{2}&BS_{1}^{T}\otimes(US_{2})^{T}\\ S_{1}B^{T}\otimes US_{2}&D_{1}\otimes\mathbbm{1}_{N_{o}}\end{pmatrix}=\Delta/\sigma_{\mathcal{R}}^{2}.$ (63)
Here, $\otimes$ denotes the Kronecker product and in the derivation, we used $\mathrm{vec}(\mathcal{A}\mathcal{B}\mathcal{C})=(\mathcal{C}^{\mathrm{T}}\otimes\mathcal{A})\mathrm{vec}(\mathcal{B})$ as well as $\mathrm{vec}(\mathcal{A}\mathcal{B})=(\mathbbm{1}_{\mathrm{dim}_{2}\mathcal{B}}\otimes\mathcal{A})\mathrm{vec}(\mathcal{B})=(\mathcal{B}^{\mathrm{T}}\otimes\mathbbm{1}_{\mathrm{dim}_{1}\mathcal{A}})\mathrm{vec}(\mathcal{A})$ where $\mathrm{dim}_{n}\mathcal{A}$ denotes the size of the $n$th dimension of $\mathcal{A}$ Petersen and Pedersen (2012). Next, we determine the eigenvalues and the rank of $\Gamma/\sigma_{x}^{2}$ and $\Delta/\sigma_{\mathcal{R}}^{2}$. After some algebra, using the determinant formula for block matrices Petersen and Pedersen (2012) and assuming $N_{i}\geq N_{h}\geq N_{o}$, we find the eigenvalues  [^6] This notation means that, e.g., in Eq. 64a, all possible combinations of $k$ and $l$ are to be taken, resulting in $N_{h}N_{o}$ distinct eigenvalues, while in Eq. 64b, each $D_{2,ll}$ occurs with multiplicity $N_{i}-N_{h}$.
[…]
$\text{multiplicity}\,N_{i}-N_{h}:$ $\quad D_{2,ll},\qquad\text{with}\quad l=1,\ldots,N_{o},$ (64b)
$\text{multiplicity}\,N_{h}N_{o}:$ $\quad 0$ (64c)
Accordingly, $\Gamma$ and $\Delta$ are positive semi-definite matrices with $\mathrm{rank}(\Gamma)=\mathrm{rank}(\Delta)=N_{o}N_{i}$. The rank deficiency of $\Gamma$ and $\Delta$ turns out to be a crucial property in the following.
III.2.4 General properties
From the results of the preceding subsection one can already infer a few general properties of the fluctuations $\delta W_{1,2}$, which become particularly simple in the case $N_{o}=1$: First, from the definition $\delta W_{1}(t)=Vz_{1}(t)=(Vz_{1,1}(t),Vz_{1,2}(t),\ldots,Vz_{1,N_{i}}(t))$, with the column vector $V$, we infer that all $N_{h}$ elements in the $i$th column of $\delta W_{1}$ fluctuate in the same way, i.e., their dynamics is controlled by the scalar $z_{1,i}(t)$, up to time-independent prefactors contained in $V$. This applies independently from the initialization. Next, we recall $\delta W_{2}(t)=z_{2}(t)A^{T}=(z_{2}(t)\cdot\bar{a}_{1},z_{2}(t)\cdot\bar{a}_{2},\ldots,z_{2}(t)\cdot\bar{a}_{N_{h}})$, where $\bar{a}_{i}$ denotes the $i$th column of $A^{T}$. In the case of zero-balanced initial conditions [described by Eq. 42 with vanishing r.h.s.], one has $1=\mathrm{rank}(W_{2}^{0})=\mathrm{rank}(W_{2}^{0,T}W_{2}^{0})=\mathrm{rank}(W_{1}^{0}W_{1}^{0,T})$ and hence $D_{1}=\mathrm{diag}(r,0,\ldots,0)$ [see Section III.2.3]. Accordingly, in this case, the scalars $z_{2,i}(t)=0$ for $i>1$, such that $\delta W_{2}(t)=(\bar{a}_{1,1},\bar{a}_{2,1},\cdots,\bar{a}_{N_{h},1})z_{2,1}(t)$, i.e., all elements of $\delta W_{2}$ fluctuate identically up to individual scaling factors determining their variance. By contrast, in the case of random initialization, $D_{1}$ is typically of maximum rank, implying that all elements of $\delta W_{2}$ fluctuate independently. Numerical results show that the above statements apply also for non-white $\Sigma^{xx}$.
III.3 Steady-state fluctuations of weights
Having obtained the linear Langevin equation [Eq. 62] for the weights in each layer, we proceed to determine the resulting steady-state statistics of the weight fluctuations. We find that, as a genuine consequence of the structure of a deep linear network, weight fluctuations are generally anisotropic. We finally provide a derivation of the IVFR.
III.3.1 Covariances
The covariance matrix $M\equiv\mathrm{cov}(\mathbf{Z})=\langle\mathbf{Z}\mathbf{Z}^{T}\rangle$ describing the fluctuations of the $z_{i}$ is determined by a Lyapunov equation Risken (1989); Laub (2005); Mandt et al. (2017); Kunin et al. (2021):
$\Gamma M+M\Gamma^{T}=\Delta.$ (65)
As one readily checks (using the fact that $\Gamma\propto\Delta$ and the identity $\mathcal{A}\mathcal{A}^{+}\mathcal{A}=\mathcal{A}$ for a matrix $\mathcal{A}$), Eq. 65 is solved by
$M\equiv\begin{pmatrix}M_{11}&M_{12}\\ M_{21}&M_{22}\end{pmatrix}=\frac{1}{2}\Gamma^{+}\Delta,$ (66)
which is symmetric of size $N_{o}(N_{i}+N_{h})\times N_{o}(N_{i}+N_{h})$, while $M_{11}\in\mathbb{R}^{N_{o}N_{i}\times N_{o}N_{i}}$, $M_{12},M_{21}^{T}\in\mathbb{R}^{N_{o}N_{i}\times N_{o}N_{h}}$, $M_{22}\in\mathbb{R}^{N_{o}N_{h}\times N_{o}N_{h}}$. This solution assumes detailed balance, which is approximately realized for SGD in the case of sufficient underparameterization [see Eq. 33 and the related discussion]. The covariances of the weight fluctuations $\delta W_{1}$ and $\delta W_{2}$ in the original space follow as
$\mathrm{cov}(\mathrm{vec}(\delta W_{1}))$ $=(\mathbbm{1}_{N_{i}}\otimes V)M_{11}(\mathbbm{1}_{N_{i}}\otimes V^{T}),$ (67a)
$\mathrm{cov}(\mathrm{vec}(\delta W_{2}))$ $=(A\otimes\mathbbm{1}_{N_{o}})M_{22}(A^{T}\otimes\mathbbm{1}_{N_{o}}),$ (67b)
with analogous relations for the cross-covariances. Since $(\mathcal{A}\otimes\mathcal{B})^{T}=\mathcal{A}^{T}\otimes B^{T}$], the eigenvalues of $M_{22}$ coincide with the eigenvalues of $\mathrm{cov}(\mathrm{vec}(\delta W_{2}))$, while the eigenvalues of $M_{11}$ coincide with the nonzero eigenvalues of $\mathrm{cov}(\mathrm{vec}(\delta W_{1}))$, noting that the latter has at least $N_{i}(N_{h}-N_{o})$ vanishing eigenvalues. A typical covariance matrix $M$ is visualized in Fig. 6. Upon averaging over the data distribution, all off-diagonal elements vanish, leaving only the diagonal ones as “universal” imprints of the data statistics (see also the discussion below). However, since we are interested in properties pertaining to a specific realization of the data, we do not perform such averaging here.
Due to the rank deficiency of $\Gamma$, the matrix $M$ is, in general, non-diagonal. This can also be inferred from the eigen-decompositions $\Gamma/\sigma_{x}^{2}=\Delta/\sigma_{\mathcal{R}}^{2}=R\Omega R^{T}$, where $\Omega$ is a diagonal matrix of rank $N_{o}N_{i}$ [see Section III.2.3], which implies $M=(R^{T})^{-1}\Omega^{+}\Omega R^{T}$ with a diagonal matrix of eigenvalues
$\Omega^{+}\Omega=\frac{\lambda\sigma_{y}^{2}}{2S}\mathbbm{1}_{N_{o}(N_{i}+N_{h})}^{N_{o}N_{i}}.$ (68)
Accordingly, $M$ is of rank $N_{o}N_{i}$ and has $N_{o}N_{h}$ vanishing eigenvalues.
Figure 6: (a) Visualization of the covariance matrix $M$ [Eq. 66] of the transformed weights $z_{i}$ for a typical simulation with $N_{i}=30$, $N_{h}=20$, $N_{o}=1$. In the special basis defined by the SVD of $W_{1}^{0}$ and $W_{2}^{0}$ [see Section III.2.3], the covariance of the second-layer weights is fully diagonal. (b,c) Eigenvalues of $M$ obtained from SGD simulations in the late-time, quasi-stationary state (symbols), compared to the theoretical predictions obtained from $M_{11}$ and $M_{22}$ in Eq. 66 (solid lines). In the simulations, we use non-white Gaussian i.i.d. samples $\mathbf{x}^{\mu}$, a label noise variance of $\sigma_{y}^{2}=0.1$, and an initial variance of the weights of (b) $\sigma_{W_{1}}^{2}=10^{-4}/N_{i}$, $\sigma_{W_{2}}^{2}=10^{-4}/N_{h}$, and (c) $\sigma_{W_{1}}^{2}=1/N_{i}$, $\sigma_{W_{2}}^{2}=1/N_{h}$. For comparison, we show the eigenvalues of $M$ obtained by assuming Gaussian i.i.d. distributed elements for $W_{1}^{0}$ and $W_{2}^{0}$ as dashed and dot-dashed lines, respectively. Eigenvalues are normalized by $\lambda\sigma_{y}^{2}/(2S)$ [cf. Eq. 70].
In order to limit complexity of the analysis, we focus in the following on a one-dimensional output $N_{o}=1$. In this case, $z_{i}$ and $\eta_{i}$ become row vectors, while $U$, $S_{2}$, $D_{2}$ become scalars (in fact, $U=\pm 1$), and
$\Gamma/\sigma_{x}^{2}=\Delta/\sigma_{\mathcal{R}}^{2}=\begin{pmatrix}D_{2}\mathbbm{1}_{N_{i}}&(US_{2})BS_{1}^{T}\\ (US_{2})S_{1}B^{T}&D_{1}\end{pmatrix}=\begin{pmatrix}V^{T}W_{2}^{0,T}W_{2}^{0}V&(W_{2}^{0}V)^{T}W_{1}^{0,T}A\\ (W_{2}^{0}V)A^{T}W_{1}^{0}&A^{T}W_{1}^{0}W_{1}^{0,T}A\end{pmatrix}.$ (69)
Using expressions for the pseudo-inverse of a block matrix Miao (1991), we find, after some tedious but straightforward algebra, that $M_{22}$ [Eq. 66] has a simple diagonal form:
$M_{22}=\mathrm{cov}(z_{2})=\frac{\lambda\sigma_{y}^{2}}{2S}\,\mathrm{diag}\left\{\frac{D_{1,ii}}{D_{1,ii}+D_{2}}\right\}_{i=1,\ldots,N_{h}}.$ (70)
Accordingly, the components of $z_{2}$ [see Section III.2.3] fluctuate independently and $M_{22,ii}$ in fact represent the eigenvalues of $\mathrm{cov}(\delta W_{2})$ [see Eq. 67b]. By contrast, $M_{11}$ is, in general, not diagonal (see Fig. 6), although its off-diagonal elements are typically small, implying that the components of $z_{1}$ are weakly cross-correlated.
III.3.2 Discussion and numerical results
We conclude this analysis with several remarks: (1) Without the cross coupling between $\delta W_{1}$ and $\delta W_{2}$ [see Section III.2.2], $\Gamma$ and $\Delta$ would be diagonal and full rank, which would imply $M\propto\mathbbm{1}_{N_{i}+N_{h}}$ and thus $[\mathrm{cov}(\delta W_{1})]\propto\mathbbm{1}_{N_{i}+N_{h}}$ and $[\mathrm{cov}(\delta W_{2})]\propto\mathbbm{1}_{N_{o}\times N_{h}}$, i.e. isotropic fluctuations. (2) In the case of Gaussian i.i.d. data considered here, the analysis starting from Section III.2.3 essentially relied only on the eigenstructure of the weight product matrices $W_{1}^{0}W_{1}^{0,T}$ and $W_{2}^{0,T}W_{2}^{0}$, which enter both the drift and the diffusion matrix [Eq. 63]. The data is assumed to be structureless [with whitened inputs, see Eq. 49] and thus appears here only as scaling factors $\sigma_{x}^{2}$, $\sigma_{y}^{2}$. Numerical results confirm that similar fluctuation spectra are also obtained for non-white data. (3) The solution in Eq. 39 only determines the $N_{\text{eff}}<N_{\text{tot}}$ degrees of freedom of $W_{2}^{0}W_{1}^{0}$ uniquely in terms of the data, while the individual weight matrices $W_{1}^{0}$ and $W_{2}^{0}$ (and hence also $D_{1,2}$) are additionally affected by the initialization. Specifically, in the limit $N_{\text{tot}}\gg N_{\text{eff}}$, the weights hardly change from their initial values during training (lazy training regime) Chizat et al. (2019); Geiger et al. (2021), allowing one to approximate a nonlinear neural network by its linearization around initialization Jacot et al. (2018); Lee et al. (2019). Assuming $W_{i}^{0}(t=0)\sim\mathcal{N}(0,\sigma_{W_{i}})$, $W_{1}^{0}W_{1}^{0,T}$ and $W_{2}^{0,T}W_{2}^{0}$ thus can be approximated as Wishart matrices, the eigenvalues of which asymptotically follow a Marchenko-Pastur distribution with characteristic magnitude $D_{i}\propto\sigma_{W_{i}}^{2}$. When initializing the weights conventionally with Glorot and Bengio (2010)
$\sigma_{W_{1}}^{2}\simeq 1/N_{i},\qquad\sigma_{W_{2}}^{2}\simeq 1/N_{h},$ (71)
and assuming $N_{h}\lesssim N_{i}$, one accordingly expects $D_{1,ii}\lesssim D_{2}$ [see Section III.2.3], such that Eq. 70 can be approximated as
$M_{22}\simeq\frac{\lambda\sigma_{y}^{2}}{2SD_{2}}D_{1},$ (72)
except for the few largest eigenvalues in $D_{1}$. We conclude that for overparameterized deep linear nets initialized in this way, weight fluctuations are generally anisotropic and their spectrum has a certain universal character encoded by the limiting eigenvalue distribution. A notable exception occurs for zero-balanced initial conditions [see Eq. 42], where, owing to $D_{1,ii}\propto\delta_{i,0}$ [see Section III.2.4], the fluctuation spectrum is isotropic. (4) The derivation leading to Eqs. 70 and 72, demonstrating the anisotropy of the fluctuation spectrum, is obtained within the large-sample number limit [$P\to\infty$; see Section III.2.1 and the related discussion], which, in the case of a single-layer net, would imply isotropic weight fluctuations [see Eq. 34]. Multi-layer nets thus induce by their structure a genuinely new effect on the fluctuation spectrum.
Figure 6(b,c) shows simulation results obtained from SGD training of a two layer linear network on (non-white) Gaussian i.i.d. data for Gaussian weight initializations [Eq. 71]] of (b) small and (c) large variance. We find good agreement with the covariance spectrum of the weights predicted by Eq. 66, which is a linearization of the full stochastic gradient flow [see Section III.2.2]. In particular, the assumption of whitened data in the theory does not appear to significantly affect the quality of the results. Confirming Eq. 70, the spectrum of the fluctuations of $W_{2}$ is essentially determined by the eigenvalues of $W_{1}^{0}W_{1}^{0,T}$ [see Section III.2.3]. Eigenvalues of $\mathrm{cov}(\delta W_{1})$ with index larger than $N_{i}$ vanish according to the linear theory [Eqs. 66 and III.3.1], while in our simulations, they are found to slowly decay towards zero with increasing index. For weights close to being zero-balanced, the fluctuations of $W_{2}$ are essentially governed by a single independent mode [see discussion in Section III.2.4], which is reflected by the prominent peak of the lowest mode in Fig. 6(b). The eigenvalue peak of the covariance of $W_{1}$ observed in (c) can be associated to an overall drift of the mean of $W_{1}$, which occurs on a large time scale. A further analysis of these and related aspects of the dynamics is reserved for a future study. Comparing panels (b) and (c), we observe that, for large values of the initial weight variance, the spectrum remains closer to the shape implied for Gaussian i.i.d. elements of $W_{1,2}^{0}$.
III.3.3 Relation to loss
Figure 7: Perturbing the loss $L$ of a two-layer linear network [Eq. 37] by the eigenvectors of the covariance matrix of the fluctuations of (a) $\delta W_{2}$ [see Eq. 73] and (b) $\delta W_{1}$. Results of training on Gaussian i.i.d. data ($\sigma_{x}^{2}=1/N_{i}$, $\sigma_{y}^{2}=1$) with SGD (connected cirlces) are compared to the theoretical predictions (solid lines), given by Eq. 73 in (a), and by the numerically evaluated $\delta L(\theta\mathbf{v}_{i})$ using the eigenvectors $\mathbf{v}_{i}$ of $\mathrm{cov}(\mathrm{vec}(\delta W_{1}))$ [Eq. 67a] in (b). The dashed-dotted line represents the prediction based on the nonlinear stochastic gradient flow. In (a), beside the exact result $\delta L=\sigma_{x}^{2}\theta^{2}D_{1,ii}/2$ (solid line), also the approximation [see Eq. 73] in terms of the eigenvalues of the covariance $M_{22}$ is plotted (dashed line). The inset shows the eigenvalues of $\mathrm{cov}(\mathrm{vec}(\delta W_{1}))$ ordered by magnitude, vs. the mode index $i$. For $i>N_{i}$, the eigenvalues abruptly drop to a smaller level, while they are predicted to vanish according to the linear model (see discussion in Section III.3.2). The corresponding drop of $\delta L$ at $i=N_{i}$ is captured by the nonlinear stochastic gradient flow simulations, while the linear theory [solid line in (b)] instead predicts a flat profile for $\delta L$. The notation $[\mathbf{v}_{i}]$ represents the reshaping of the eigenvector $\mathbf{v}_{i}$ into a matrix compatible with $W_{1}$. We used $N_{o}=1$, $N_{h}=40$, $N_{i}=50$ in (a) and $N_{o}=1$, $N_{h}=20$, $N_{i}=30$ in (b). The behaviors for whitened and non-whitened data are similar.
We now study how the fluctuation variance relates to the loss, thereby providing an analytical derivation of the IVFR Feng and Tu (2021). The reduced quadratic loss [see Section III.2.2] is given by $\tilde{L}(z_{1},z_{2})=\frac{1}{2}\sigma_{x}^{2}\mathrm{tr}[D_{2}z_{1}z_{1}^{T}+z_{2}^{T}z_{2}D_{1}+z_{2}^{T}(US_{2})z_{1}(BS_{1}^{T})]$, which indicates that the perturbation of the loss due to fluctuations of $z_{1}$ and $z_{2}$ is controlled by the spectrum of $W_{2}^{0,T}W_{2}^{0}$ and $W_{1}^{0}W_{1}^{0,T}$, respectively [see Section III.2.3]. In order to capture the effect of the fluctuations in the original weight space, we start from Eq. 37 and recall that the principal modes $\mathbf{a}_{i=1,\ldots,N_{h}}$ of the fluctuations of $W_{2}$ (i.e., the columns of $A$) fulfill $W_{1}^{0}W_{1}^{0,T}\mathbf{a}_{i}=D_{1,ii}\mathbf{a}_{i}$ with $D_{1,ii}\simeq 2SD_{2}M_{22,ii}/(\lambda\sigma_{y}^{2})$ [see Sections III.3.1 and 72]. Using the basic relation $\mathrm{tr}(\mathcal{A}\mathcal{B})=\mathrm{tr}(\mathcal{B}\mathcal{A})$ as well as the fact that $W_{2}^{0}W_{1}^{0}=\sigma_{x}^{-2}\Sigma_{0}^{yx}$ for whitened inputs, the perturbation of the loss [Eq. 37] along $\mathbf{a}_{i}$ follows, after a bit of algebra, as [cf. Eq. 35]:
$\delta L(\theta\mathbf{a}_{i})=L(W_{1}^{0},W_{2}^{0}+\theta\mathbf{a}_{i}^{T})-L(W_{1}^{0},W_{2}^{0})=\frac{1}{2}\sigma_{x}^{2}\theta^{2}D_{1,ii}\simeq\frac{SD_{2}\sigma_{x}^{2}}{\lambda\sigma_{y}^{2}}\theta^{2}M_{22,ii}\,.$ (73)
[…]
By contrast, for a perturbation of the loss along eigendirections $\mathbf{v}_{i}$ of $\mathrm{cov}(\mathrm{vec}\,\delta W_{1})$ [Eq. 67a], we do not observe any dependence on the mode index $i$, i.e., $\delta L(\theta\mathbf{v}_{i})=L(W_{1}^{0}+\theta[\mathbf{v}_{i}],W_{2}^{0})-L(W_{1}^{0},W_{2}^{0})\simeq\mathrm{const}.$, even when the weight covariance is anisotropic [see Fig. 7(b)]. Here, $[\mathbf{v}_{i}]$ represents the vector $\mathbf{v}_{i}$ reshaped into the form of $W_{1}^{0}$. Note, however, that $\delta L$ determined within the linearized fluctuation theory [i.e., using Eq. 66] does not capture the bimodal structure of $\delta L$ observed in numerical solutions of SGD or SGF [Section III.2.1], but instead remains flat. We thus conclude that, for a two-layer linear network, the IVFR is generally not fulfilled for perturbations of the first-layer weights $W_{1}$.
III.4 Relaxation dynamics
We illustrate here how the linearized SGF derived in the previous section can be used to analyze the layer-wise relaxation dynamics of the weights. Focusing on $N_{o}=1$, we diagonalize the (symmetric) drift and diffusion matrices as $\Gamma/\sigma_{x}^{2}=\Delta/\sigma_{\mathcal{R}}^{2}=R\Omega R^{T}$ [see Eq. 69], with a diagonal $\Omega=\mathrm{diag}(\omega_{k})$ ($k=1,\ldots,N_{h}+N_{i})$ and an orthogonal matrix $R$ with eigenvectors in the columns. According to Section III.2.3, the eigenvalues are given by
$\omega_{k=1,\ldots,N_{h}}=D_{1,kk}+D_{2},\qquad\omega_{k=N_{h}+1,\ldots,N_{i}}=D_{2},\qquad\omega_{k=N_{i}+1,\ldots,N_{i}+N_{h}}=0,$ (74)
with $D_{1,2}$ defined in Section III.2.3. Note that all $\omega_{k}\geq 0$. Introducing new dynamical variables $\bm{\chi}=R^{T}\mathbf{Z}$ and noises $\bm{\zeta}=R^{T}\bm{\eta}$, we transform Eq. 62 to
$\dot{\chi}_{k}(t)=\sigma_{x}^{2}\omega_{k}\chi_{k}(t)+\zeta_{k}(t),\qquad\langle\zeta_{k}(t)\zeta_{k^{\prime}}(t^{\prime})\rangle=\sigma_{\mathcal{R}}^{2}\,\omega_{k}\,\delta_{kk^{\prime}}\,\delta(t-t^{\prime}).$ (75)
The general solution is given by $\chi_{k}(t)=e^{-\omega_{k}(t-t_{0})}\chi_{k}(t_{0})+\int_{t_{0}}^{t}\mathrm{d}s\,e^{\omega_{k}(s-t)}\zeta_{k}(s)$ with some initial time $t_{0}$. Accordingly, in the stationary state, one obtains the correlation functions
$\langle\chi_{k}(t)\chi_{k^{\prime}}(0)\rangle$ $=\langle\chi_{k}(0)\chi_{k^{\prime}}(0)\rangle e^{-\sigma_{x}^{2}\omega_{k}t},$ (76)
$\langle\chi_{k}(t)\chi_{k^{\prime}}(t)\rangle$ $=e^{-\sigma_{x}^{2}(\omega_{k}+\omega_{k^{\prime}})t}\langle\chi_{k}(0)\chi_{k^{\prime}}(0)\rangle+\delta_{kk^{\prime}}\frac{\sigma_{\mathcal{R}}^{2}}{2\sigma_{x}^{2}}\left(1-e^{-2\sigma_{x}^{2}\omega_{k}t}\right),$ (77)
where we assumed time-translation invariance, which allows us to chose the arbitrary time origin as $t_{0}=0$. Modes $\chi_{k}$ belonging to eigenvalues $\omega_{k}=0$ remain fixed at their initialization $\chi_{k}(t)=\chi_{k}(0)$. By contrast, modes with $\omega_{k}>0$ relax towards zero [Eq. 76], while their variance saturates at late times at the equilibrium value [Eq. 77]:
$\langle\chi_{k}\chi_{k^{\prime}}\rangle_{\text{eq}}=\begin{cases}\langle\chi_{k}(0)\chi_{k^{\prime}}(0)\rangle\qquad&\omega_{k}=0=\omega_{k^{\prime}}\\ \delta_{kk^{\prime}}\frac{\sigma_{\mathcal{R}}^{2}}{2\sigma_{x}^{2}}\qquad&\text{otherwise}.\\ \end{cases}$ (78)
Notably, as a consequence of the equality of $\Gamma$ and $\Delta$ (up to an overall scale), the fluctuations of the modes $\chi$ are isotropic (independent of $k$) at late times and only determined by the relative strength of the noise ($\sigma_{\mathcal{R}}^{2}$) and the loss ($\sigma_{x}^{2}$). Since the initial values of $W_{1,2}$ are uncorrelated, we may assume $\langle\chi_{k}(0)\chi_{k^{\prime}}(0)\rangle\propto\delta_{kk^{\prime}}$. In order to transform back into the original weight space, one uses $\delta W_{1}=V[\mathcal{P}_{1}R\bm{\chi}]^{T}$ and $\delta W_{2}=[\mathcal{P}_{2}R\bm{\chi}]^{T}A^{T}$ [see Section III.2.3 and notes that $\mathbf{Z}$ is a column vector], with the projection matrices $\mathcal{P}_{1}=\begin{pmatrix}\mathbbm{1}_{N_{i}}&0\\ 0&0\end{pmatrix}$ and $\mathcal{P}_{2}=\begin{pmatrix}0&0\\ 0&\mathbbm{1}_{N_{h}}\end{pmatrix}$, both of size $(N_{i}+N_{h})^{2}$. From these considerations we infer that the relaxation dynamics of the weights is in general a superposition of exponentials, dominated by the smallest eigenvalue $\omega_{k}$. A more detailed investigation of the noise-driven time evolution is reserved for a future work.
IV Summary
In the present study, we have analytically and numerically investigated the fluctuating dynamics of linear networks with mean-square loss in the late-time, stationary state of training. The considered single-layer linear network trained with SGD solves the usual least-squares regression problem. Besides being of theoretical interest, linear models describe the stationary training regime of deep nonlinear networks (see Appendix A) and form the foundation to construct SGD-based Bayesian-sampling methods (see Appendix B). Two-layer linear networks share important properties with deep nonlinear nets, such as nonlinear training dynamics, a non-convex loss landscape, and a strong dependence on initial conditions Saxe et al. (2014); Laurent and von Brecht (2018); Arora et al. (2019). We have considered Gaussian i.i.d. training data with targets generated by linearly transforming the inputs and adding noise (including the limit of fully random targets). We have numerically demonstrated that our results hold also for correlated data. Since steady-state fluctuations vanish for overparameterized nets in our training setup, we have focused on the underparameterized case, where sample number is larger than the input (or data) dimension. Our main motivation was to study the simplest possible case that gives rise to the inverse-variance-flatness relation (IVFR) Feng and Tu (2021) and to gain analytical insights into the nature of the SGD noise and the induced weight fluctuations.
We have obtained the following main results:
  1. 1.
    We have shown that one source of anisotropy of the weight fluctuations is the broken detailed balance of SGD, which emerges in the presence of label noise for mildly underparameterized nets, while it disappears in the infinite sample limit. In the latter case, the noise covariance becomes equal to the Hessian of the loss.
  2. 2.
    For two-linear nets, we have demonstrated that another source of anisotropy stems from the rank deficiency of the drift and diffusion matrices that describe the stochastic gradient flow of the weight fluctuations in each layer. This is a generic consequence of the structure of a deep linear network and is thus expected to arise also for nonlinear networks.
  3. 3.
    Based on these insights, we have provided an analytical derivation of the IVFR for a two-layer linear network from first principles. Our approach therefore complements previous analyses of this relation Feng and Tu (2021); Yang et al. (2023) and suggest that the IVFR generally arises for networks with at least two layers.
  4. 4.
    As a byproduct of our investigation, we have characterized the layer-wise continuum dynamics of SGD in two-layer linear nets, extending previous works that considered the deterministic gradient flow Saxe et al. (2014); Lampinen and Ganguli (2019); Tarmoun et al. (2021); Braun et al. (2022).
Understanding weight fluctuations in neural networks is not only interesting from a theoretical perspective Engel and Van den Broeck (2001); Bahri et al. (2020); Huang (2021), but has a number of practical implications. In fact, noise-induced transition events are known to be relevant for optimization in a rough loss landscape Zhu et al. (2019); Daneshmand et al. (2018); Nguyen et al. (2019); Mori et al. (2022). Recently, the fluctuation-dissipation ratio, which has its origin in statistical physics Cugliandolo (2011); Marconi et al. (2008), has been introduced as a novel training measure Yaida (2019). In the context of differential privacy Abadi et al. (2016), random noise is added to the gradient updates and interacts with the SGD noise. It will be interesting to extend and further characterize the mathematical properties of the solutions obtained here, which – despite the linearity of the considered network models – involve some nontrivial aspects of random matrix theory. For a future work, we also envisage a detailed analysis of more sophisticated neural network models, including ones with larger depths and nonlinear activation functions.
Appendix A Nonlinear networks
The linear approximation applies to a nonlinear network in the stationary state of training, provided fluctuations are sufficiently small, which requires large batch sizes and small learning rates (see the discussion in Section II.2.2). While nonlinear networks also reduce to linear models in the large-width limit Jacot et al. (2018); Lee et al. (2019), steady-state weight fluctuations are absent in the overparameterized case [see Eq. 16]. However, provided the mean dynamics is sufficiently slow, we expect that a linearization is viable also in the transient training phase.
We denote the output function of the nonlinear network as $f(\mathbf{x},\mathbf{w})$, where $\mathbf{w}\in\mathbb{R}^{N}$ collectively represents the network parameters. Taylor expansion around a reference point $\mathbf{w}^{*}$ renders
$f(\mathbf{x},\mathbf{w})\simeq f(\mathbf{x},\mathbf{w}^{*})+(\mathbf{w}-\mathbf{w}^{*})\cdot\nabla_{\mathbf{w}}f(\mathbf{x},\mathbf{w}^{*})=b(\mathbf{x})+\mathbf{w}\cdot\nabla_{\mathbf{w}}f(\mathbf{x},\mathbf{w}^{*}),$ (79)
with a bias term $b(\mathbf{x})=f(\mathbf{x},\mathbf{w}^{*})-\mathbf{w}^{*}\cdot\nabla_{\mathbf{w}}f(\mathbf{x},\mathbf{w}^{*})$ that is independent of $\mathbf{w}$. Defining a Jacobian matrix
$J=\begin{pmatrix}\nabla_{\mathbf{w}}^{T}f(\mathbf{x}_{1},\mathbf{w}^{*})\\ \vdots\\ \nabla_{\mathbf{w}}^{T}f(\mathbf{x}^{P},\mathbf{w}^{*})\end{pmatrix}\in\mathbb{R}^{P\times N},$ (80)
we can write the MSE loss as
$L(\mathbf{w})=\frac{1}{2P}\sum_{\mu}[f(\mathbf{x}^{\mu},\mathbf{w})-y^{\mu}]^{2}=\frac{1}{2P}||\mathbf{w}^{T}J-Y||^{2}+\mathrm{const}.$ (81)
up to a constant independent of $\mathbf{w}$. Accordingly, the linear framework of Section II can be adopted by replacing the design matrix $X$ by the Jacobian matrix $J$ [see Eq. 3], which effectively renders a random feature model in terms of the features $\partial_{w_{i}}f(\mathbf{x},\mathbf{w}^{*})$ Cao et al. (2019); Wei et al. (2022).
Appendix B Connection to Bayesian methods
Since Bayesian neural networks involve the weight distribution as a central ingredient, it is informative to briefly review this approach here Wilson and Izmailov (2020); Jospin et al. (2022). The prediction $y^{*}$ for a new data point $x^{*}$ is computed as an average of the likelihood $p(y^{*}|x^{*},w)$ weighted by the posterior $p(w|\mathcal{D})$: $p(y^{*}|x^{*},\mathcal{D})=\int_{w}p(y^{*}|x^{*},w)p(w|\mathcal{D})\mathrm{d}w$. According to Bayes’ rule, the posterior distribution is given by $p(w|\mathcal{D})\propto p(y|x,w)p(w)=\exp(-\mathcal{L}_{\text{reg}}(x,y,w))$, where $\mathcal{D}=\{x^{\mu},y^{\mu}\}_{\mu}$ denotes the training data and $p(w)$ the prior distribution of the weights, while $p(y|x,w)\propto\exp(-\mathcal{L}(x,y,w))$ is the likelihood function defined in term of the loss $\mathcal{L}$. The regularized loss is given accordingly as $\mathcal{L}_{\text{reg}}=-\log p(w|\mathcal{D})=\mathcal{L}(x,y,w)+\chi||w||^{2}_{2}$, where the last equation follows by assuming a Gaussian prior, which gives rise to an $L_{2}$-regularization. Classical (non-stochastic) training amounts to approximating the posterior as $p(w|\mathcal{D})=\delta(w-w_{\text{map}})$ using the point estimate $w_{\text{map}}=\text{argmin}_{w}\mathbb{E}_{\mathcal{D}}\mathcal{L}_{\text{reg}}(x,y,w)$, where $\mathcal{L}_{\text{reg}}(x,y,w)=-\log p(w|\mathcal{D})=\mathcal{L}(x,y,w)-\log p(w)$ and $\mathbb{E}_{\mathcal{D}}$ denotes a batch average over the training data. In the stationary state of SGD, instead, one samples from Eqs. 30 and 32, which describe a different distribution from the true posterior. In order to use SGD as a Bayesian sampler, several methods have been proposed in the literature, such as preconditioning Li et al. (2015); Mandt et al. (2017) or injecting Gaussian white noise into the SGD dynamics Welling and Teh (2011); Ma et al. (2015); Adhikari et al. (2023). This isotropic white noise is constructed such that it dominates over the minibatch-sampling noise in the limit of small learning rates, inducing $C+Q\propto\mathbbm{1}$ in Eq. 30.
Appendix C Covariance of mini-batches
Consider (possibly vectorial) quantities
$G_{B}=\frac{1}{S}\sum_{\mu\in\mathcal{B}}g^{\mu},\qquad G=\langle G_{B}\rangle_{\mathcal{B}}=\frac{1}{P}\sum_{\mu=1}^{P}g^{\mu},$ (82)
where $G_{B}$ is defined as an average over mini-batches $\mathcal{B}$ of size $S=|\mathcal{B}|$, while $G$ is given by the average of $G_{B}$ over the full batch of size $P>S$. The covariance of $G_{B}$ und mini-batch sampling with replacement can be calculated, following similar steps as in Appendix E of Ziyin et al. (2022):
$\begin{split}\mathrm{cov}(G_{B})&\equiv\langle(G_{B}-\langle G_{B}\rangle_{\mathcal{B}})(G_{B}-\langle G_{B}\rangle_{\mathcal{B}})^{T}\rangle_{\mathcal{B}}=\langle G_{B}G_{B}^{T}\rangle_{\mathcal{B}}-GG^{T}\\ &=\frac{1}{S}\left[\frac{1}{P}\sum_{\mu=1}^{P}g^{\mu}(g^{\mu})^{T}-GG^{T}\right].\end{split}$ (83)
The corresponding expression for sampling without replacement takes the same form but with the prefactor $(P-S)/(S(P-1))$ instead of $1/S$ Ziyin et al. (2022). The two cases coincide if $S=1$ or $P\gg S$.
Appendix D Additive noise approximation
We justify in the following the use of the additive noise approximation in the stationary training regime of a linear network. We start from the Langevin Eq. 13, which we write
$\dot{\mathbf{w}}=-H(\mathbf{w}-\mathbf{w}^{*})+\sqrt{\eta}\bm{\zeta},\qquad\eta\equiv\frac{\lambda}{S}$ (84)
with $\mathbf{w}^{*}$ given by Eq. 7. The correlations of the noise $\bm{\zeta}$ are described by the large-sample number limit [Eq. 19]:
$\langle\bm{\zeta}(t)\bm{\zeta}(t^{\prime})^{T}\rangle=SC(\mathbf{w})\delta(t-t^{\prime})=\left[H\bar{\mathbf{w}}\bar{\mathbf{w}}^{T}H+\mathrm{tr}(H\bar{\mathbf{w}}\bar{\mathbf{w}}^{T})H+\sigma_{\epsilon}^{2}H\right]\delta(t-t^{\prime}),$ (85)
which is sufficient for the purpose of estimating the importance of the various contributions. We consider $\eta$ as a smallness parameter and assume an expansion of the form
$\mathbf{w}=\mathbf{w}_{0}+\sqrt{\eta}\mathbf{w}_{1}+\mathcal{O}(\eta).$ (86)
Analogously, we expand the noise as $\bm{\zeta}=\bm{\zeta}_{0}+\bm{\zeta}_{1}+\ldots$, where the correlations of the new noises $\bm{\zeta}_{i}$ follow successively by evaluating Eq. 85 with Eq. 86.
Inserting Eq. 86 into Eq. 84 readily yields
$\mathbf{w}_{0}=\mathbf{w}^{*}.$ (87)
Next, using $\mathbf{w}^{*}=\mathbf{u}+H^{-1}J$ [Eq. 7] renders
$\langle\bm{\zeta}_{0}(t)\bm{\zeta}_{0}(t^{\prime})^{T}\rangle=C_{0}\delta(t-t^{\prime}),\qquad C_{0}=[JJ^{T}+\mathrm{tr}(JJ^{T}H^{-1})H+\sigma_{\epsilon}^{2}H]\sim\sigma_{x}^{2}\sigma_{\epsilon}^{2},$ (88)
and the Langevin equation at the order $\sqrt{\eta}$:
$\dot{\mathbf{w}}_{1}=-H\mathbf{w}_{1}+\bm{\zeta}_{0}.$ (89)
The variance of the resulting stationary solution is given by [c.f. Eq. 33]
$\langle\mathbf{w}_{1}\mathbf{w}_{1}^{T}\rangle=\frac{1}{2}H^{-1}C_{0}\sim\sigma_{\epsilon}^{2}$ (90)
Using this result in the expansion of Eq. 85 provides the $\mathcal{O}(\eta)$-term of the noise correlation:
$\langle\bm{\zeta}_{1}(t)\bm{\zeta}_{1}(t^{\prime})^{T}\rangle=C_{1}\delta(t-t^{\prime}),\qquad C_{1}=\frac{\eta}{2}\left[H\langle w_{1}w_{1}^{T}\rangle H+\mathrm{tr}(H\langle w_{1}w_{1}^{T}\rangle)H\right]\sim\eta\sigma_{x}^{4}\sigma_{\epsilon}^{2}$ (91)
and a corresponding $\mathcal{O}(\eta^{1/2})$-term with correlation $C_{1/2}=\sqrt{\eta/2}\,J\mathbf{w}_{1}^{T}H+\ldots\sim\sqrt{\eta}\sigma_{x}^{2}\sigma_{\epsilon}^{2}$ stemming from the cross terms between $\mathbf{w}^{*}$ and $\mathbf{w}_{1}$. A comparison of the relative magnitude of these contributions and the one in Eq. 88 indicates that the multiplicative noise $\bm{\zeta}_{1}$ remains subdominant in the stationary state if $\lambda/S\ll 1$ and $\lambda\sigma_{x}^{2}/S\ll 1$ are fulfilled, which is the case for typical setups. A more rigorous analysis based on adiabatic elimination techniques Venturelli and Gross (2022); Gardiner (2009) for the Fokker-Planck equation [Eq. 28] is reserved for a future work.
Appendix E Noise covariance
Assuming whitened inputs, Sections III.2.2 and 56 imply:
$\lambda^{-1}\mathrm{cov}(\mathcal{R}_{1})_{ij,kl}$ $=(W_{2}^{0,T}\Sigma_{0}^{yx})_{il}(\Sigma_{0}^{yx,T}W_{2}^{0})_{jk}+\sigma_{x}^{2}\langle y^{2}\rangle\delta_{jl}(W_{2}^{0,T}W_{2}^{0})_{ik}-\sigma_{x}^{2}(W_{2}^{0,T}\Sigma_{0}^{yx}w^{*,T}W_{2}^{0})_{ik}\delta_{jl}$
$\qquad-\sigma_{x}^{2}(W_{2}^{0,T}w^{*}\Sigma_{0}^{yx,T}W_{2}^{0})_{ik}\delta_{jl}-\sigma_{x}^{2}(W_{2}^{0,T}\Sigma_{0}^{yx})_{il}(w^{*,T}W_{2}^{0})_{jk}-\sigma_{x}^{2}(W_{2}^{0,T}w^{*})_{il}(\Sigma_{0}^{yx,T}W_{2}^{0})_{jk}$
$\qquad+\sigma_{x}^{4}(W_{2}^{0,T}w^{*}w^{*,T}W_{2}^{0})_{ik}\delta_{jl}+\sigma_{x}^{4}(W_{2}^{0,T}w^{*})_{il}(w^{*,T}W_{2}^{0})_{jk}$
$=\sigma_{x}^{2}\langle y^{2}\rangle\delta_{jl}(W_{2}^{0,T}W_{2}^{0})_{ik}-(W_{2}^{0,T}\Sigma_{0}^{yx}\Sigma_{0}^{yx,T}W_{2}^{0})_{ik}\delta_{jl}=\sigma_{x}^{2}\langle y^{2}\rangle\delta_{jl}\left(1-1/P\right)(W_{2}^{0,T}W_{2}^{0})_{ik},$ (92)
$\lambda^{-1}\mathrm{cov}(\mathcal{R}_{2})_{ij,kl}$ $=(\Sigma_{0}^{yx}W_{1}^{0,T})_{il}(W_{1}^{0}\Sigma_{0}^{yx,T})_{jk}+\sigma_{x}^{2}\langle y^{2}\rangle\delta_{ik}(W_{1}^{0}W_{1}^{0,T})_{jl}-\sigma_{x}^{2}(\Sigma_{0}^{yx}W_{1}^{0,T})_{il}(W_{1}^{0}w^{*,T})_{jk}$
$\qquad-\sigma_{x}^{2}(\Sigma_{0}^{yx}w^{*,T})_{ik}(W_{1}^{0}W_{1}^{0,T})_{jl}-\sigma_{x}^{2}(w^{*}W_{1}^{0,T})_{il}(W_{1}^{0}\Sigma_{0}^{yx,T})_{jk}-\sigma_{x}^{2}(w^{*}\Sigma_{0}^{yx,T})_{ik}(W_{1}^{0}W_{1}^{0,T})_{jl}$
$\qquad+\sigma_{x}^{4}(w^{*}W_{1}^{0,T})_{il}(W_{1}^{0}w^{*,T})_{jk}+\sigma_{x}^{4}(w^{*}w^{*,T})_{ik}(W_{1}^{0}W_{1}^{0,T})_{jl}$
$=\sigma_{x}^{2}\langle y^{2}\rangle\delta_{ik}\left(1-1/P\right)(W_{1}^{0}W_{1}^{0,T})_{jl},$ (93)
$\lambda^{-1}\mathrm{cov}(\mathcal{R}_{1},\mathcal{R}_{2})_{ij,kl}$ $=(W_{2}^{0,T}\Sigma_{0}^{yx}W_{1}^{0,T})_{il}\Sigma_{0,jk}^{yx,T}+\sigma_{x}^{2}\langle y^{2}\rangle W^{0,T}_{2,ik}W^{0,T}_{1,jl}-\sigma_{x}^{2}(W_{2}^{0,T}\Sigma_{0}^{yx}w^{*,T})_{ik}W_{1,jl}^{0,T}$
$\qquad-\sigma_{x}^{2}(W_{2}^{0,T}w^{*}\Sigma_{0}^{yx,T})_{ik}W_{1,jl}^{0,T}-\sigma_{x}^{2}(W_{2}^{0,T}\Sigma_{0}^{yx}W_{1}^{0,T})_{il}w^{*,T}_{jk}-\sigma_{x}^{2}(W_{2}^{0,T}w^{*}W_{1}^{0,T})_{il}\Sigma_{0,jk}^{yx,T}$
$\qquad+\sigma_{x}^{4}(W_{2}^{0,T}w^{*}w^{*,T})_{ik}W_{1,jl}^{0,T}+\sigma_{x}^{4}(W_{2}^{0,T}w^{*}W_{1}^{0,T})_{il}w^{*,T}_{jk}$
$=\sigma_{x}^{2}\langle y^{2}\rangle W^{0,T}_{2,ik}W^{0,T}_{1,jl}-(W_{2}^{0,T}\Sigma_{0}^{yx}\Sigma_{0}^{yx,T})_{ik}W_{1,jl}^{0,T}=\sigma_{x}^{2}\langle y^{2}\rangle(1-1/P)W^{0,T}_{2,ik}W^{0,T}_{1,jl},$ (94)
which leads to Section III.2.2.